### v3

# Deep Computerized Adaptive Testing

Abstract
Computerized adaptive tests (CATs) play a crucial role in educational assessment and diagnostic screening in behavioral health. Unlike traditional linear tests that administer a fixed set of pre-assembled items, CATs adaptively tailor the test to an examinee’s latent trait level by selecting a smaller subset of items based on their previous responses. Existing CAT applications predominantly rely on item response theory (IRT) models with a single latent variable, a choice driven by both conceptual simplicity and computational feasibility. However, many real-world item response datasets exhibit complex, multi-factor structures, limiting the applicability of CATs in broader settings. In this work, we develop a novel CAT system that incorporates multivariate latent traits, building on recent advances in Bayesian sparse multivariate IRT [doi:10.1080/01621459.2025.2476786]. Our approach leverages direct sampling from the latent factor posterior distributions, significantly accelerating existing information-theoretic item selection criteria by eliminating the need for computationally intensive Markov Chain Monte Carlo (MCMC) simulations. Recognizing the potential suboptimality of existing item selection rules which are often based on myopic one-step-lookahead optimization of some information-theoretic criterion, we propose a double deep Q-learning algorithm to learn an optimal item selection policy. Through simulation and real-data studies, we demonstrate that our approach not only accelerates existing item selection methods but also highlights the potential of reinforcement learning in CATs. Notably, our Q-learning-based strategy consistently achieves the fastest posterior variance reduction, leading to earlier test termination. In finite-horizon settings, it also yields final posterior distributions that more closely approximate the oracle posterior distribution—the posterior that would be obtained if test takers had responded to all items in the item bank.
Keywords: Reinforcement Learning; Multidimensional Item Response Theory; Adaptive Data Collection; Deep Q-Learning;
1 Introduction
Multidimensional Computerized Adaptive Testing (MCAT) has revolutionized the field of educational and psychological assessments by dynamically selecting tailored items from a large test pool, thereby enhancing the efficiency and precision of latent ability estimates [van2010elements]. Powered by Multidimensional Item Response Theory (MIRT) [bock2021item], MCAT leverages multidimensional statistical inference to evaluate respondents’ multidimensional latent traits, and allows for a more comprehensive and efficient assessments compared to unidimensional approaches. MCAT’s adaptability and accuracy are also particularly crucial in high-stakes diagnostic assessments such as in clinical psychology and psychiatry, where it can be substituted for in-person assessments by clinical professionals especially in areas with limited medical resources [gibbons2016computerized].
A substantial body of research in MCAT has focused on item selection strategies derived from the experimental design principles. One prominent strategy is to select items that maximize the determinant of the Fisher information matrix evaluated at the current estimates of the latent traits [segall1996multidimensional, segall2000principles], known as the D-optimality criterion. An alternative, the A-optimality criterion, aims to minimize the trace of the asymptotic covariance matrix, thereby reducing overall estimation variance [a584ed95-55e4-36a9-8d67-ff8dd3f06250]. In the absence of nuisance latent abilities, both A-optimality and D-optimality have demonstrated superior accuracy relative to other common experimental design criteria [mulder2009multidimensional].
Another prominent line of MCAT item selection algorithms leverages Kullback-Leibler (KL) information, often within a Bayesian framework. A common approach is to select items that produce response distributions at the true latent trait value, $\theta_{0}$, that differs maximally from the response distributions generated at the other value of $\theta$ [Chang1996AGI, pekl]. Moving beyond response distributions alone, some researchers propose maximizing the KL divergence between the current posterior distribution and the posterior distribution at the next selection step, thereby enhancing adaptation through updated trait estimates [kl-is]. Another promising strategy is the mutual information criterion, which aims to maximize entropy reduction of the current posterior distribution, encouraging more and more accurate posterior estimates [mi_selection]. In particular, [selection_survey] demonstrates both theoretical and empirical advantages of the Bayesian mutual information item selection rule over common experimental criteria such as D-optimality. More detailed theoretical comparison of KL information and Fisher information criteria is presented by [Wang_Chang_Boughton_2011].
Although numerous effective item selection rules have been proposed in the MCAT literature, they all rely on one-step lookahead optimization of an information-theoretic criterion. Despite their ease of implementation and attractive theoretical properties, these selection rules are inherently myopic: they select items based solely on immediate information gain, ignoring how current choices influence future decisions, which can lead to suboptimal policies. For example, existing methods tend to favor items with high discrimination parameters [Chang2015]. However, CAT researchers also recommend reserving highly discriminating items for the later stages of testing to improve efficiency [doi:10.1177/01466219922031338]. Integrating heuristic guidance into existing selection rules still remains challenging.
Addressing these limitations, we propose a novel deep CAT system that integrates a flexible Bayesian MIRT model with a non-myopic online item selection policy, guided by reinforcement learning (RL) principles [Sutton1998]. Leveraging recent advancements in Bayesian sparse MIRT [doi:10.1080/01621459.2025.2476786], our framework seamlessly accommodates multiple latent factors with complex loading structures, while maintaining scalability in both the number of items and factors. To learn the optimal item selection policy prioritizing the assessment of target factors, we draw on contemporary RL methodologies and introduce a general double deep Q-learning algorithm [Mnih2015, van_Hasselt_Guez_Silver_2016]. This algorithm efficiently trains a Q-network offline using only item parameter estimates, which can be subsequently deployed online to select optimal items based on the current multivariate latent factor posterior distribution. When the test terminates, our framework robustly characterizes the full latent factor posterior distributions rather than providing only a point estimate.
A primary contribution of our work is to show how identification of the latent-factor posterior distribution leads to substantial computational gains in online item selection. Given that such posterior distribution is deemed to be non-Gaussian and unknown, traditional Bayesian methods typically rely on expensive MCMC simulations [BeguinGlas2001] and combined with additional data augmentation to handle categorical likelihood [f3928584-48a1-3ab1-b404-2ec9ffdf51bb, Polson01122013]. Our approach achieves substantial acceleration by directly sampling latent factor posterior distributions [doi:10.1080/01621459.2025.2476786, Durante_2019, Botev_2016]. Notably, this improvement not only increases the efficiency of existing Bayesian item-selection procedures but also provides a computational foundation for training our proposed Q-learning algorithm through rapid, large-scale simulations of testing sessions.
Another critical advancement in our work is the integration of CAT within a reinforcement learning framework. This approach addresses the practical necessity of prioritizing accurate estimation in high-dimensional settings and overcomes known limitations of greedy item selection methods in sequential decision-making [10.5555/560669]. Building on the Bayesian MIRT foundation, our neural-network architecture incorporates the identified posterior parameters as state variables, allowing the model to learn optimal item-selection policies through large amount of testing simulations. This formulation bridges the two paradigms: the Bayesian component provides statistically grounded representations of examinee uncertainty, while the reinforcement-learning component leverages these representations to optimize sequential decisions. The trained neural network is deployable on standard laptops without GPU acceleration, making it suitable for online adaptive testing applications. The sequential nature of CAT aligns naturally with deep Q-learning methods, which have demonstrated remarkable success across diverse application domains [Silver2016, Kalashnikov2018QTOptSD]. Notably, RL has been successfully employed in educational measurement settings to design personalized learning plans and enhance learning outcomes [doi:10.3102/10769986221129847, doi:10.1177/0146621619858674].
Finally, our work aligns with the emerging research trend of framing traditional statistical sequential decision-making problems as optimal policy learning tasks [modern_bad_review]. This perspective has appeared in Bayesian adaptive design [bad_review, bad, Foster2021DeepAD] and Bayesian optimization [srinivas2010gaussian, NIPS2016_5ea1649a], where recent methods aim to move beyond one-step criteria toward policy learning that account for long-term consequences. Our contribution fits within this broader trend by providing a principled approach to cognitive and behavioral assessment.
The remainder of the paper is organized as follows. Section 2 motivates our deep CAT framework using a cognitively complex item bank from a recent clinical study [cat-cog]. Section 3 reviews existing information-theoretic item selection methods and outlines the necessity to reformulating CAT as a reinforcement learning task. Section 4 introduces a general Bayesian framework that accelerates existing CAT item-selection rules and serves as the foundation for our Q-learning algorithm. Section 5 details our reinforcement learning approach, including the neural network architecture and double Q-learning algorithm. Finally, Sections 6 and 7 evaluates our method through both simulations and real data experiments. The implementations of the baseline Bayesian CAT algorithms, our Q-learning algorithm, and the driver codes to reproduce all the experiments in the paper can be found in https://github.com/JiguangLi/deep_CAT.
2 Adaptive High-Dimensional Cognitive Assessment
Our proposed deep CAT system is motivated by the growing need for adaptive cognitive assessment in high-dimensional latent spaces. Cognitive impairment, particularly Alzheimer’s dementia (AD), is a major public health challenge, affecting 6.9 million individuals in the U.S. in 2024 [alzheimers2024facts]. Early detection of mild cognitive impairment (MCI), a precursor to AD, is crucial for slowing disease progression and improving patient outcomes [10.1001/jamanetworkopen.2018.1726]. However, traditional neuropsychological assessments are costly, time-consuming, and impractical for frequent use, underscoring the need for more efficient and scalable assessment methods.
Recently, pCAT-COG, the first computerized adaptive test (CAT) item bank based on MIRT for cognitive assessment, demonstrated its potential as an alternative to clinician-administered evaluations [cat-cog]. The data were collected from $730$ participants. After careful item calibration and model selections, the final item bank consisted of $J=57$ items covering five cognitive subdomains: episodic memory, working memory, executive function, semantic memory, and processing speed. Since each item comprised three related tasks, we used a binary score, where $1$ indicates correct answers on all tasks for our analysis.
Following [cat-cog], we fit a six-factor bifactor model [bifactor] to the $730$ by $57$ binary response matrix, with one primary factor representing the global cognitive ability and five secondary factors. This yields a $57$ by $6$ factor loading matrix, visualized in Figure 1, where rows correspond to pCAT-COG items and columns represent distinct factors.
Figure 1: Estimated Bifactor Factor Loading Matrix for pCAT-COG
Since items in pCAT-COG are primarily designed to measure the general cognitive factor and only partially capture subdomain information, our goal is to develop an item selection strategy that efficiently estimates the primary factor (first column) using as few items as possible while maintaining robustness to subdomain influences. Additionally, the selection algorithm must be computationally efficient to navigate a six-dimensional latent space in real time for seamless interactive testing. Online item selection is critical, as learning an optimal sequence from $57$ items offline presents an intractable combinatorial problem, even in a binary response setting.
Adaptive cognitive assessment is essential for cognitive assessment, as designing items is costly and administering all items is time-consuming. Our deep CAT system requires on average $11.2$ items to reduce the posterior variance of the primary factor from $1$ to $0.16$ (posterior s.d. =0.4=0.4), whereas the next-best mutual information method requires $12.6$ items to achieve the same precision. The item bank is currently expanding to $500$ items with a more nationally representative sample. Success on this prototype dataset paves the way for broader deployment in clinical research.
3 From One-Step Optimization to Policy Learning
We formally define the problem of designing a CAT system from a reinforcement learning (RL) perspective. Typical CAT systems consist of two components:
  • Offline Calibration: An MIRT model is fitted to a calibrated dataset $Y\in\mathbbm{R}^{N\times J}$ to estimate the item characteristic parameters, where $N$ represents the number of examinees, and $J$ represents the number of items in the item bank.
  • Online Deployment: Given the estimated item parameters, an item selection algorithm is deployed online to adaptively select items for future examinees.
The performance of CAT can be measured by the number of items required to estimate an examinee’s latent traits with sufficient precision.
3.1 Notations and Problem Setups
For the entirety of the paper, we assume the calibration dataset $Y$ is binary, where each element $y_{ij}$ represents whether subject $i$ answered item $j$ correctly. We consider a general two-parameter MIRT framework with $K$ latent factors [bock2021item]. Let $B\in\mathbbm{R}^{J\times K}$ denote the factor loading matrix containing the discrimination parameters, and $D\in\mathbbm{R}^{J}$ denote the intercept vector. For each examinee $i$ with multivariate latent trait $\bm{\bm{\theta}}^{(i)}\in\mathbbm{R}^{K}$, the data-generating process for $y_{ij}$ is given by:
$y_{ij}|\bm{\bm{\theta}}^{(i)},B_{j},D_{j}\sim\text{Bernoulli}(\Phi(B_{j}^{\prime}\bm{\bm{\theta}}^{(i)}+D_{j})),$ (1)
where $B_{j}$ is the $j$-th row of $B$, $D_{j}$ is the $j$-th entry of $D$, and $\Phi(.)$ denotes the standard normal cumulative distribution function. The item parameters can be compactly expressed as $\{\bm{\xi}_{j}\}_{j=1}^{J}:=\{(B_{j},D_{j})\}_{j=1}^{J}$. This two-parameter MIRT framework is highly general, imposing no structural constraints on the loading matrix and requiring no specific estimation algorithms for item parameter calibration.
Given the estimated item parameters, we need to design an item selection algorithm that efficiently tests a future examinee with unobserved latent ability $\bm{\bm{\theta}}\in\mathbbm{R}^{K}$. The sequential nature of the CAT problem makes Bayesian approaches particularly appealing. Without loss of generality, we assume a standard multivariate Gaussian prior on $\bm{\bm{\theta}}\sim\mathcal{N}(0,\mathbbm{I}_{K})$, and define the following notations:
  • For any positive integer $n$, let $[n]$ denote all the positive integers no greater than $n$. Let $j\in[J]$ be the index for the items in the item bank.
  • Let $j_{t}$ denote the index of the item selected at step $t$ based on an arbitrary item selection rule. Define $\mathcal{I}_{t}$ as the set of the first $t$ administered items, and let $R_{t}:=[J]\setminus\mathcal{I}_{t-1}$, the set of available items before the $t$-th item is picked.
  • We use the shorthand notation $f(\bm{\bm{\theta}}|Y_{1:T})$ to represent the latent factor posterior distributions $f(\bm{\bm{\theta}}|Y_{1:T},\bm{\xi}_{1:T})$ after $T$ items have been selected. Here, the response history is denoted by $Y_{1:T}:=[y_{j_{1}},\cdots,y_{j_{T}}]^{\prime}$, and the item parameters are given by $\bm{\xi}_{1:T}:=(B_{1:T},D_{1:T})$, where $B_{1:T}:=[B_{j_{1}},\cdots,B_{j_{T}}]^{\prime}$ and $D_{1:T}:=[D_{j_{1}},\cdots,D_{j_{t}}]^{\prime}$.
At time $(t+1)$, the algorithm takes the current posterior $f(\bm{\bm{\theta}}|Y_{1:t})$ as input and outputs the next item selection $j_{t+1}$. This process continues until at time $T^{\prime}$, either when the posterior variance of $f(\bm{\bm{\theta}}|Y_{1:T^{\prime}})$ falls below a predefined threshold $\tau^{2}$, or when $T^{\prime}=H$, where $H\leq J$ is the maximum number of items that can be administered. Hence the goal of CAT is to minimize $T^{\prime}$ given the estimated item parameters of the item bank.
3.2 Reviews of KL Information Item Selection Rules
Given that our proposed deep CAT system is built on a general Bayesian MIRT framework [doi:10.1080/01621459.2025.2476786], we briefly revisit common Bayesian item selection rules and show in Section 4.1 how our framework can be used to accelerate these baseline methods. The popular KL Expected A Priori (EAP) rule selects the $t$-th item based on the average KL information between response distributions on the candidate item at the EAP estimate $\hat{\bm{\bm{\theta}}}_{t-1}=\int_{\bm{\bm{\theta}}}\bm{\bm{\theta}}f(\bm{\bm{\theta}}|Y_{1:(t-1)})d\bm{\bm{\theta}}$, and random factor $\bm{\bm{\theta}}$ sampled from the posterior distribution $f(\bm{\bm{\theta}}|Y_{1:(t-1)})$ [pekl]:
$\arg\max_{j_{t}\in R_{t}}\int_{\bm{\bm{\theta}}}\bigg\{\sum_{l=0}^{1}P(y_{j_{t}}=l|\hat{\bm{\bm{\theta}}}_{t-1})\log\frac{P(y_{j_{t}}=l|\hat{\bm{\bm{\theta}}}_{t-1})}{P(y_{j_{t}}=l|\bm{\bm{\theta}})}\bigg\}f(\bm{\bm{\theta}}|Y_{1:(t-1)})d\bm{\bm{\theta}}.$ (2)
Rather than focusing on the KL information based on the response distributions, [kl-is] propose the MAX Pos approach by maximizing the KL information between two subsequent latent factor posteriors $f(\bm{\theta}|Y_{1:(t-1)})$ and $f(\bm{\theta}|Y_{t})$. Intuitively, this approach prioritizes items that induce the largest shift in the posterior, formalized as:
$\arg\max_{j_{t}\in R_{t}}\sum_{y_{j_{t}}=0}^{1}f(y_{j_{t}}|Y_{1:(t-1)})\int_{\bm{\theta}}f(\bm{\theta}|Y_{1:(t-1)})\log\frac{f(\bm{\theta}|Y_{1:(t-1)})}{f(\bm{\theta}|Y_{1:t})}d\bm{\theta},$ (3)
where $f(y_{j_{t}}|Y_{1:(t-1)})$ represents the posterior predictive probability:
$f(y_{j_{t}}|Y_{1:(t-1)})=\int_{\bm{\theta}}f(y_{j_{t}}|\bm{\theta})f(\bm{\theta}|Y_{1:(t-1)})d\bm{\theta}.$ (4)
A third Bayesian strategy, the mutual information (MI) approach, maximizes the mutual information between the current posterior distribution and the response distribution of new item $y_{j_{t}}$ [mi_selection]. Rooted in experimental design theory [Rnyi1961OnMO], we can also interpret mutual information as the entropy reduction of the current posterior estimation of $\bm{\theta}$ after observing new response $y_{j_{t}}$. More formally:
$\arg\max_{j_{t}\in R_{t}}I_{M}(\bm{\theta},y_{j_{t}})=\arg\max_{j_{t}\in R_{t}}\sum_{y_{j_{t}}=0}^{1}\int_{\bm{\theta}}f(\bm{\theta},y_{j_{t}}|Y_{1:(t-1)})\log\frac{f(\bm{\theta},y_{j_{t}}|Y_{1:(t-1)})}{f(\bm{\theta}|Y_{1:(t-1)})f(y_{j_{t}}|Y_{1:(t-1)})}d\bm{\theta}.$ (5)
Given the impressive empirical success for the MI method [selection_survey], We also introduce a competitive heuristic item-selection rule that chooses items with the highest predictive variance under the current posterior estimates, serving as an additional benchmark method. Formally, write $c_{j_{t}}:=f(y_{j_{t}}|Y_{1:(t-1)})$ as defined in (4), and consider the following criterion:
$\text{argmax}_{j_{t}\in R_{t}}\int_{\bm{\theta}}(\Phi(B_{j_{t}}^{\prime}\bm{\theta}+D_{j_{t}})-c_{j_{t}})^{2}f(\bm{\theta}|Y_{1:(t-1)})d\bm{\theta}.\\$ (6)
This rule selects the item $j_{t}$ that maximizes the variance of the predictive means, weighted by the current posterior $f(\bm{\theta}|Y_{1:(t-1)})$. In Appendix A, we establish a connection between this selection rule and the MI method, showing that both favor items with high prediction uncertainty.
3.3 Reinforcement Learning Formulation
Existing item selection rules face several limitations: first, these methods rely on one-step-lookahead selection, choosing items that provide the most immediate information without considering their impact on future selections, which can result in suboptimal policies [Sutton1998]. Second, they are heuristically designed to balance information across all latent factors, rather than emphasizing the most essential factors of interest. In the pCAT-COG study, the goal is to accurately measure the global cognition factor, rather than estimating all latent factors with equal precision. Finally, these rules do not directly minimize test length, potentially increasing test duration without proportional gains in accuracy.
A more principled approach is to formulate item selection as a reinforcement learning (RL) problem, where the optimal policy is learned using Bellman optimality principles rather than relying on heuristics that ignore long-term planning. Beyond addressing myopia, an RL-based formulation provides a direct mechanism to minimize the number of items required to reach a predefined posterior variance reduction threshold, explicitly guide item selection toward accurately measuring the primary factors of interests.
More generally, consider a general finite horizon setting where each examinee can answer at most $H\leq J$ items, and the CAT algorithm terminates whenever the posterior variance of all factors of interest is smaller than a certain threshold $\tau^{2}$, or when $H$ is reached. Since $H$ can be set sufficiently large, it serves as a practical secondary stopping criterion to prevent excessively long tests. Formally:
  • State Space $\mathcal{S}$: the space of all possible latent factor posteriors $\mathcal{S}:=\{f(\bm{\theta}|Y_{1:t}):t\subset\{1,\cdots,H\}\}.$ The state variable in CAT represents the Bayesian estimate of the examinee’s latent traits as a multivariate distribution at time $t$. We may write $s_{t}:=f(\bm{\theta}|Y_{1:t})$, with $s_{0}:=\mathcal{N}(0,\mathbbm{I}_{k})$. The state space is discrete and can be exponentially large. Since there are $J$ items in the item bank and the responses are binary, the total number of possible states is $\sum_{t=1}^{H}\binom{J}{t}\cdot 2^{t}$.
  • Action space $\mathcal{A}$: the item bank $J]:=\{1,\cdots,J\}$. However, at the t-th selection time, the action space is the remaining items in the test bank that have not yet been selected $R_{t}:=[J]\setminus\mathcal{I}_{t-1}$, as we do not select the same item twice.
  • Transition Kernel $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$, where $\Delta(\mathcal{S})$ is the state of probability distribution over $\mathcal{S}$. The transition kernel $\mathcal{P}$ specifies the stochastic rule that governs how the state evolves after an action is taken. In the CAT context, the kernel maps the current latent-trait estimate $s_{t}$ and the selected item $a_{t}$ to the next posterior state $s_{t+1}$. Conditional on $(s_{t},a_{t})$, the next estimate $s_{t+1}$ depends on the examinee’s binary response $y_{t}\in\{0,1\}$ to item $a_{t}$, which yields only two possible posterior updates. Because the true ability of the examinee and thus the response probability $p(y_{t}=1\mid s_{t},a_{t})$ are unknown, the transition kernel in CAT is also unknown and must be approximated.
  • Reward Function: To directly minimize the test length, we assign a simple 010-1 reward structure, where we assign negative rewards whenever more items are needed to reduce posterior variance to a given threshold. In a $K$-factor MIRT model, we often prioritize a subset of factors $\mathcal{K}\subset[K]$ (e.g. $\mathcal{K}=\{1\}$ for pCAT-COG), with the test terminating once the posterior variances of all factors in $\mathcal{K}$ fall below the predefined threshold $\tau^{2}$:
    $R^{(t)}(s_{t},a,s_{t+1})=\begin{cases}-1&\text{if }V_{t+1}>\tau^{2},\\ 0&\text{otherwise},\end{cases}$ (7)
    where $V_{t+1}=\max_{k\in\mathcal{K}}\text{Var}(\bm{\theta}_{k}|Y_{t})$ is the maximum marginal posterior variance among the prioritized factors $\mathcal{K}$. This also simplifies learning the value function, as the rewards are always bounded integers within [H,1][-H,-1]. A similar reward has been successfully deployed in adaptive learning settings [doi:10.3102/10769986221129847].
Rather than following a heuristic rule, we learn a policy $\pi$ that maps the current state (latent factor estimates) to a distribution over potential items:
$\pi:\ \mathcal{S}\to\Delta(\mathcal{A}),\qquad a_{t}\sim\pi(\cdot\mid s_{t}).$
Given an initial state $S_{0}=s_{0}$, and the discount factor $\gamma\in(0,1]$, the value function is
$v_{\pi}(s_{0}):=\mathbbm{E}_{\pi}\left[\sum_{t=0}^{H-1}\gamma^{t}R^{(t)}(s_{t},\pi(s_{t}),s_{t+1})|S_{0}=s_{0}\right],$ (8)
The expectation $\mathbb{E}_{\pi}[\cdot]$ is taken over trajectories generated by drawing $a_{t}$ from $\pi(\cdot\mid s_{t})$ and then $s_{t+1}$ from the transition kernel $\mathcal{P}(\cdot\mid s_{t},a_{t})$. Because direct evaluation of (8) is challenging, it is more convenient to consider the action–value function
$Q_{\pi}(s_{0},a):=\mathbbm{E}_{\pi}\left[\sum_{t=0}^{H-1}\gamma^{t}R^{(t)}(s_{t},\pi(s_{t}),s_{t+1})|S_{0}=s_{0},A_{0}=a\right],$ (9)
with $v_{\pi}(s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}Q_{\pi}(s,a)$. The optimal policy $\pi^{\star}$ satisfies the Bellman equation [10.5555/560669]:
$Q_{\pi^{\star}}(s,a)\;=\;\mathbb{E}_{S^{\prime}\sim\mathcal{P}(\cdot\mid s,a)}\!\left[\,R(s,a,s^{\prime})+\gamma\,\max_{a^{\prime}}Q_{\pi^{\star}}(s^{\prime},a^{\prime})\,\right],$ (10)
where $s^{\prime}$ is the next posterior estimate after applying action $a$ to the current posterior $s$, and $v_{\pi^{\star}(s)}=\max_{a}Q_{\pi^{\star}}(s,a)$. Since the state space grows exponentially and the transition kernel is unknown, solving for $\pi^{*}$ using traditional dynamic programming approaches become intractable [Sutton1998].
In practice, our deep Q-learning approach [Mnih2013PlayingAW] does not evaluate these expectations analytically. Let the transition $(s,a,r,s^{\prime})$ represent one testing step, where item $a$ is selected under state $s$, the reward $r$ is observed, and the posterior is updated to $s^{\prime}$ through the Bayesian MIRT model. We approximate the Bellman fixed point in (10) by fitting a parametric function $Q_{w}$ that minimizes the squared temporal-difference loss:
$\mathcal{L}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\big[\big(Q_{w}(s,a)-y(s,a,r,s^{\prime})\big)^{2}\big]$ (11)
where
$y(s,a,r,s^{\prime})=r+\gamma\max_{a^{\prime}}Q_{\bar{w}}(s^{\prime},a^{\prime}).$ (12)
Here, $\mathcal{D}$ (replay buffer) denotes a collection of previously observed testing steps $(s,a,r,s^{\prime})$ generated during simulation, which is used to approximate the expectation above by empirical averaging. The parameter $\bar{w}$ corresponds to a delayed copy of the model parameters $w$, updated less frequently to stabilize the numerical optimization. We describe the full Q-learning algorithm in Section 5.
4 Accelerating Item Selection via Posterior Identification
The central insight of our deep CAT framework is that, by iteratively applying the E-step of the PXL-EM algorithm [doi:10.1080/01621459.2025.2476786], the latent factor posteriors can be shown to have tractable posterior updates. This result is critical for reinforcement learning, as the posterior distribution is deemed to be non-Gaussian and is analytically intractable under the traditional MIRT literature. By obtaining a tractable representation of this posterior, we can parametrize the examinee’s evolving ability and uncertainty as a well-defined state variable, which is an essential prerequisite for applying Q-learning to adaptive testing. The Bayesian MIRT formulation thus not only replaces costly MCMC procedures with efficient posterior updates under a probit link for existing Bayesian item selection rules discussed in Section 3.2, but also supplies the statistical foundation that makes the subsequent reinforcement-learning framework feasible.
Specifically, we show that the latent factor posterior updates during CAT belong to an instance of the unified-skew-normal distribution [unified_skew_normal], defined as follows:
Definition 4.1.
Let $\Phi_{T}\left\{V;\Sigma\right\}$ represent the cumulative distribution functions of a T-dimensional multivariate Gaussian distribution $N_{T}\left(0_{T},\Sigma\right)$ evaluated at vector $V$. A K-dimensional random vector $\bm{\theta}\sim\operatorname{SUN}_{K,T}(\mu,\Omega,\Delta,\gamma,\Gamma)$ has the unified skew-normal distribution if it has the probability density function:
$\phi_{K}(\bm{\theta};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mu},\Omega)\frac{\Phi_{T}\left\{\gamma+\Delta^{\prime}\bar{\Omega}^{-1}\omega^{-1}(\bm{\theta}-\mu);\Gamma-\Delta^{\prime}\bar{\Omega}^{-1}\Delta\right\}}{\Phi_{T}(\gamma;\Gamma)}.$
Here, $\phi_{K}(\bm{\theta};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mu},\Omega)$ is the density of a $K$-dimensional multivariate Gaussian with expectation $\mu=\left(\mu_{1},\ldots,\mu_{K}\right)^{\prime}$, and a $K$ by $K$ covariance matrix $\Omega=\omega\bar{\Omega}\omega$, where $\bar{\Omega}$ is the correlation matrix and $\omega$ is a diagonal matrix with the square roots of the diagonal elements of $\Omega$ in its diagonal. $\Delta$ is a $K$ by $T$ matrix that determines the skewness of the distribution, and $\gamma\in\mathbbm{R}^{T}$ control the flexibility in departures from normality.
In addition, the $(K+T)\times(K+T)$ matrix $\Omega^{*}$, having blocks $\Omega_{[11]}^{*}=\Gamma,\Omega_{[22]}^{*}=\bar{\Omega}$ and $\Omega_{[21]}^{*}=\Omega_{[12]}^{*^{\prime}}=\Delta$, needs to be a full-rank correlation matrix.
Suppose an arbitrary CAT item selection algorithm has already selected $T$ items with item parameters $\bm{\xi}_{1:T}=(B_{1:T},D_{1:T})$, where $B_{1:T}:=[B_{j_{1}},\cdots,B_{j_{T}}]^{\prime}$ and $D_{1:T}:=[D_{j_{1}},\cdots,D_{j_{t}}]^{\prime}$. Then it is possible to show the following result:
Theorem 4.2.
Consider a K-factor CAT item selection procedure after selecting $T$ items, with $\mathcal{N}(0,\mathbbm{I}_{K})$ prior placed on the test taker’s latent trait $\bm{\theta}$. If $Y_{1:T}=\left(y_{j_{1}},\ldots,y_{j_{T}}\right)^{\prime}$ is conditionally independent binary response data from the two-parameter probit MIRT model definde in (1), then
$(\bm{\theta}\mid Y_{1:T},B_{1:T},D_{1:T})\sim\operatorname{SUN}_{K,T}\left(\mu_{\mathrm{post}},\Omega_{\text{post }},\Delta_{\text{post }},\gamma_{\text{post }},\Gamma_{\text{post }}\right),$
with posterior parameters
$\mu_{\mathrm{post}}=0_{K},\quad\Omega_{\text{post }}=\mathbbm{I}_{K},\quad\Delta_{\text{post }}=C_{1}^{\prime}C_{3}^{-1},$
$\gamma_{\text{post }}=C_{3}^{-1}C_{2},\quad\Gamma_{\text{post }}=C_{3}^{-1}\left(C_{1}C_{1}^{\prime}+\mathbbm{I}_{T}\right)C_{3}^{-1},$
where $C_{1}=\text{diag}(2y_{j_{1}}-1,\cdots,2y_{j_{T}}-1)B_{1:T}$ and $C_{2}=\text{diag}(2y_{j_{1}}-1,\cdots,2y_{j_{T}}-1)D_{1:T}$. The matrix $C_{3}$ is a $T$ by $T$ diagonal matrix, where the $(t,t)$-th entry is $(\|B_{t,T}\|_{2}^{2}+1)^{\frac{1}{2}}$, with $B_{t,T}$ representing the $t$-th row of $B_{1:T}$.
Theorem 4.2 should be read as a fully Bayesian finite sample statement. With a multivariate normal prior on the latent trait $\bm{\theta}$ and a probit MIRT likelihood, the posterior $f(\bm{\theta}\mid Y_{1:n})$ belongs to the unified skew normal family. This does not conflict with the empirical Bayes result of [ChangStout1993], which establishes asymptotic posterior normality as the number of items $J$ grows. Theorem 4.2 therefore provides the exact finite sample form of $p(\bm{\theta}\mid Y_{1:n})$, which is especially relevant for CAT, where the number of administered items at any time is small and large sample normal approximations are unreliable. This characterization enables exact posterior calculations without large sample approximations, which improves the fidelity of uncertainty quantification and the resulting item selection policy in short tests.
According to [unified_skew_normal], the distribution of an arbitrary unified skew normal distribution $\bm{\theta}\sim\operatorname{SUN}_{K,T}(\mu,\Omega,\Delta,\gamma,\Gamma)$ has a stochastic representation as a linear combination of a $K$-dimensional multivariate normal random variable $V_{0}$, and a $T$-dimensional truncated multivariate normal random variable $V_{1,-\gamma}$ as follows:
$\bm{\theta}\overset{d}{=}\mu+\omega(V_{0}+\Delta\Gamma^{-1}V_{1,-\gamma}),$ (13)
where $V_{0}\sim\mathcal{N}(0,\bar{\Omega}-\Delta\Gamma^{-1}\Delta^{\prime})\in\mathbbm{R}^{K}$, and $V_{1,-\gamma}$ is a multivariate normal distribution $\mathcal{N}(0,\Gamma)$ truncated to $\{V_{1}\in\mathbb{R}^{T}:V_{1i}\geq-\gamma_{i}\quad,\forall i\}$.
Based on Theorem 4.2, we have a closed-form expression for the posterior parameters $(\mu,\Omega,\Delta,\gamma,\Gamma)$ for $\theta$. Recall that $\Omega=\omega\hat{\Omega}\omega$ is the standard covariance correlation matrix decomposition from definition 4.1, and hence the only unrealized stochastic terms in equation (13) are $V_{0}$ and $V_{1,-\gamma}$. This suggests sampling from the latent factor posterior distribution $f(\bm{\theta}|Y_{1:T})$ requires two independent steps: sampling from a $K$-dimensional multivariate normal distribution $V_{0}$, and a $T$-dimensional multivariate truncated normal distribution $V_{1}$. As a result, the direct sampling approach scales efficiently with the number of factors $K$, since generating samples from $K$-dimensional multivariate normal distributions is trivial. Moreover, in CAT settings, the test is typically terminated early, meaning $T$ remains relatively small. When $T$ is moderate (e.g, $T<1,000$), sampling from the truncated multivariate normal distribution remains computationally efficient using the minimax tilting method [Botev_2016]. The exact proof of Theorem 4.2 and the sampling details can be found in Appendix B.
The direct sampling approach provides substantial gains in both computational efficiency and numerical precision compared with traditional MCMC methods commonly used in the MIRT literature [BeguinGlas2001, JiangTemplin2019]. In standard MIRT setting, the posterior distribution of the latent factors is typically regarded as intractable and non-Gaussian, requiring a Markov chain to be constructed via data-augmentation techniques [f3928584-48a1-3ab1-b404-2ec9ffdf51bb, Polson01122013] so that its stationary distribution approximates the posterior. This procedure entails repeated simulation of augmented data and is inherently sequential, which limits opportunities for parallelization. Moreover, it demands additional tuning, burn-in, and convergence diagnostics to ensure that the chain adequately converges to the posterior distribution. In contrast, Theorem 4.2 establishes that the stationary distribution can be expressed exactly as a unified skew-normal distribution, allowing direct and parallel sampling without the need for iterative convergence procedures.
Theorem 4.2 also plays a central role in our proposed deep Q-learning algorithm, as it fully characterizes the latent factor posterior distribution $f(\bm{\theta}|Y_{1:T})$, enabling parametrization of the state variable $s_{T}$, which serves as an input to the Q-network illustrated in Section 5.1. We also illustrate the application of Theorem 4.2 in accelerating the existing item selection rules below.
4.1 Accelerating Existing Rules
While Bayesian CAT criteria require multidimensional integration, these quantities are typically evaluated in practice via Monte Carlo approximation rather than analytic quadrature in moderate to high dimensions. In such Bayesian CAT implementations, the dominant computational cost lies in obtaining valid samples from the latent factor posterior. By providing exact posterior samples without Markov chain construction, our approach removes this primary bottleneck and renders the remaining Monte Carlo integration computationally efficient and easily parallelizable.. For example, in computing the KL-EAP item selection rule (2), we directly sample from $f(\bm{\theta}|Y_{1:(t-1)})$ instead of resorting to MCMC, and evaluate the integral via Monte-Carlo integration. Since the EAP estimate $\hat{\bm{\theta}}_{t-1}$ remains fixed at time step $t$, evaluating the KL information term is straightforward. For the Max Pos item selection rule in equation (3), we again obtain i.i.d. samples from $f(\bm{\theta}|Y_{1:(t-1)})$ and compute the posterior predictive probabilities in (4). Although the density ratio $\frac{f(\bm{\theta}|Y_{1:(t-1)})}{f(\bm{\theta}|Y_{1:t})}$ is difficult to evaluate, we can leverage the conditional independence assumption of the MIRT model. In particular, we can express the joint distribution of $(\bm{\theta},y_{j_{t}})$ as
$f(\bm{\theta},y_{j_{t}}|Y_{1:(t-1)})=f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}})f(y_{j_{t}}|Y_{1:(t-1)})=f(y_{j_{t}}|\bm{\theta})f(\bm{\theta}|Y_{1:(t-1)}).$ (14)
Using Equation (14), we rewrite the KL information term in Equation (3) as:
$\int f(\bm{\theta}|Y_{1:(t-1)})\log\frac{f(\bm{\theta}|Y_{1:(t-1)})}{f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}})}d\bm{\theta}$ $=\int f(\bm{\theta}|Y_{1:(t-1)})\log\frac{f(\bm{\theta}|Y_{1:(t-1)})f(y_{j_{t}}|Y_{1:(t-1)})}{f(\bm{\theta},Y_{j_{t}}|Y_{1:(t-1)})}d\bm{\theta}$
$=\int f(\bm{\theta}|Y_{1:(t-1)})\log\frac{f(y_{j_{t}}|Y_{1:(t-1)})}{f(y_{j_{t}}|\bm{\theta})}d\bm{\theta}.$
Since $f(y_{j_{t}}|\bm{\theta})$ can be easily computed for each $\bm{\theta}\sim f(\bm{\theta}|Y_{1:(t-1)})$, the online-computation of Mas Pos remains efficient.
Although the mutual information selection rule in Equation (5) has demonstrated strong empirical performance [selection_survey], its computational complexity remains a significant challenge. By applying Equation (14), we can rewrite mutual information as
$\arg\max_{j_{t}\in R_{t}}\sum_{y_{j_{t}}=0}^{1}f(y_{j_{t}}|Y_{1:(t-1)})\int_{\bm{\theta}}f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}})\log\frac{f(y_{j_{t}}|\bm{\theta})}{f(y_{j_{t}}|Y_{1:(t-1)})}d\bm{\theta}.$ (15)
This formulation reveals that maximizing mutual information is structurally similar to Max Pos but requires much more computational effort. Unlike Max Pos, where sampling is performed from the current posterior $f(\bm{\theta}|Y_{1:(t-1)})$, mutual information requires sampling from future posteriors $f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}})$. Since each candidate item $j_{t}\in R_{t}$ has two possible outcomes ($y_{j_{t}}=0$ or $y_{j_{t}}=1$), evaluating equation (15) requires obtaining samples from $|R_{t}|\times 2$ distinct posterior distributions. Even if sampling each individual posterior is computationally efficient, this approach becomes impractical for large item banks.
We hence propose a new approach to dramatically accelerate the computation of the mutual information quantity using the idea of importance sampling and resampling and bootstrap filter [Gordon1993NovelAT, sir]. Rather than explicitly sampling the future posterior $f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}})$ for each $j_{t}\in R_{t}$ and $y_{j_{t}}\in\{0,1\}$, we can simply sample from the current posterior $f(\bm{\theta}|Y_{1:(t-1)})$ once, and then perform proper posterior reweighting. Under equation (1), let $p(\bm{\theta})$ be the prior on the latent factors $\bm{\theta}$, $l_{1}(\bm{\theta})$ be the current data likelihood, and $l_{2}(\bm{\theta})$ denote the future data likelihood after observing $y_{j_{t}}$. We have the future posterior density as below:
$f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}})$ $\propto l_{2}(\bm{\theta})p(\bm{\theta})\propto\frac{l_{2}(\bm{\theta})p(\bm{\theta})}{l_{1}(\bm{\theta})p(\bm{\theta})}f(\bm{\theta}|Y_{1:(t-1)})$
$=\Phi(B_{j_{t}}^{\prime}\bm{\theta}+D_{j_{t}})^{y_{j_{t}}}(1-\Phi(B_{j_{t}}^{\prime}\bm{\theta}+D_{j_{t}}))^{1-y_{j_{t}}}f(\bm{\theta}|Y_{1:(t-1)}).$ (16)
Equation (4.1) suggests we can generate samples from $f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}})$ via reweighting and resampling from the current posterior samples $f(\bm{\theta}|Y_{1:(t-1)})$. Specifically, given a sufficient large set of posterior samples $(\bm{\theta}_{1},\cdots,\bm{\theta}_{M})\sim f(\bm{\theta}|Y_{1:(t-1)})$, we assign distinct weights for each sample $m\in\{1,\cdots,M\}$ and future item $j_{t}\in R_{t}$:
$q_{m}=\frac{w_{m}}{\sum_{i=1}^{M}w_{i}},\quad\text{where }w_{i}=\Phi(B_{j_{t}}^{\prime}\bm{\theta}_{i}+D_{j_{t}})^{y_{j_{t}}}(1-\Phi(B_{j_{t}}^{\prime}\bm{\theta}_{i}+D_{j_{t}}))^{1-y_{j_{t}}}.$
To sample from $f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}})$, we then draw from the discrete distribution over $(\bm{\theta}_{1},\cdots,\bm{\theta}_{M})$, placing weight $q_{m}$ on $\bm{\theta}_{m}$. This approach eliminates the need to sample from $|R_{t}|\times 2$ distinct posteriors directly, further accelerating mutual information computation and making it scalable for large item banks.
5 Learning Optimal Item Selection Policy
Building on the Bayesian MIRT foundation established in the previous section, we now formulate the problem of CAT as an RL task. The Bayesian framework provides two key advantages that make this integration both principled and computationally feasible. First, the identified latent factor posterior distributions offer a well-defined representation of examinee knowledge, allowing their corresponding posterior parameters to be parametrized directly as state variables. Without such identification, encoding an unknown and analytically intractable posterior would be ambiguous. Second, because reinforcement learning typically requires extensive simulations to learn an optimal policy, the acceleration achieved in online item selection enables rapid simulation of testing sessions, thereby substantially improving the efficiency of policy training.
As illustrated in Section 3.3, an RL approach addresses the myopic nature of traditional CAT selection rules, enables a more flexible reward structure, and directly minimizes the number of items required for performing online adaptive testing. Specifically, we propose a novel double Q-learning algorithm [Mnih2015, van_Hasselt_Guez_Silver_2016] for online item selection in CAT. The algorithm trains a deep neural network offline using only the item parameters estimated from an arbitrary two-parameter MIRT model. During online item selection, the neural network takes the current posterior distribution $s_{t}:=f(\bm{\theta}\mid Y_{1:t})$ as input and outputs the next item selection for step $(t+1)$. The neural network architecture is described in Section 5.1, while the double deep Q-learning algorithm is detailed in Section 5.2.
5.1 Deep Q-learning Neural Network Design
A potential reason why a deep reinforcement learning approach has not been proposed in the CAT literature is due to the ambiguity arising from unidentified latent factor distributions. However, by Theorem 4.2, it is straightforward to compactly parametrize the posterior distribution $s_{t}:=f(\bm{\theta}|Y_{1:t})$ at each time step $t$ using the parameters $C_{1}\in\mathbbm{R}^{t\times K}$, $C_{2}\in\mathbbm{R}^{t}$, and $\text{diag}(C_{3})\in\mathbbm{R}^{t}$ , where $\text{diag}(C_{3})$ represents the diagonal vector of $C_{3}$. This structured representation of the posterior enables item selection policy learning via deep neural networks, which takes posterior parameters as input and outputs item selection.
Neural networks can be regarded as flexible function approximators that learn nonlinear mappings between inputs and outputs through a composition of simple transformations [Goodfellow-et-al-2016]. In our framework, the deep neural network consists of two key components: an encoder and a classifier. At each time step $t$, the encoder maps the collection of posterior parameters to a latent representation in $\mathbb{R}^{L}$, where $L$ behaves a hyperparameter and represents the dimension of the latent feature space. The classifier then takes this latent representation and predicts the next item from the item bank that is expected to yield the highest reward.
Define the collection of the posterior parameters $\tilde{\bm{\xi}}_{t}:=\{(C_{1h},C_{2h},C_{3h})\}_{h=1}^{t}$ , where $C_{1h}\in\mathbb{R}^{K}$ is the $h$-th row of $C_{1}$, $C_{2h}\in\mathbb{R}$ is the $h$-th element of $C_{2}$, and $C_{3h}\in\mathbb{R}$ is the h-th elements in the diagonal of $C_{3}$. Since the size of the posterior parameters $\tilde{\bm{\xi}}_{t}$ grows over time, and permuting the tuples within $\tilde{\bm{\xi}}_{t}$ still describes the same posterior, we have to design a neural network that can take inputs of growing size, and can provide output that is permutation invariant of the inputs. One solution is to consider the weight sharing idea from the Bayesian experimental design literature [Foster2021DeepAD]. Let $\phi_{1}(.):\mathbb{R}^{K+2}\rightarrow\mathbb{R}^{L_{1}}$ denote an encoder component that maps each tuple $(C_{1h},C_{2h},C_{3h})\in\mathbb{R}^{K+2}$ to an $L_{1}$-dimensional latent space, and consider the operation $g_{1}(.)$ as follows:
$g_{1}(\tilde{\bm{\xi}}_{t}):=\sum_{h=1}^{t}\phi_{1}\{(C_{1h},C_{2h},C_{3h})\}.$ (17)
Observe that $g_{1}(.)$ is capable of handling growing number of inputs through summations of $\phi_{1}(.)$ functions over $t$. More importantly, permuting the order of the tuples in $\tilde{\bm{\xi}}_{t}$ does not change the value of $g_{1}(.)$, since summation is permutation invariant. Although the form of $g_{1}$ may look restrictive, any function $f(.)$ operating on a countable set can be decomposed into the form $\rho\circ g_{1}(.)$, where $\rho(.)$ is a suitable transformation that can be learned from another neural network (see Theorem 2 of [10.5555/3294996.3295098]).
In principle, if we define the classification component of the neural network as $\rho(.):\mathbb{R}^{L_{1}}\rightarrow\mathbb{R}^{J}$, the network design $\rho\circ g_{1}(.)$, coupled with the Q-learning algorithm is sufficient to learn the optimal policy, and can be deployed offline with significantly less computational cost since it does not require either sampling or optimization. However, to enhance learning efficiency, we enrich the input space $\tilde{\bm{\xi}}_{t}$ with a matrix of prediction quartiles $\Psi_{t}\in\mathbbm{R}^{J\times Q}$, where each row contains a vector of quantiles of the predictive distribution for item $j$. We form these quantiles by first drawing samples $\bm{\theta}_{1},\cdots,\bm{\theta}_{M}\sim f(\bm{\theta}|Y_{1:t})$ as in Section 4, and then computing the quantiles of the prediction samples $\{\Phi(B_{j}^{\prime}\bm{\theta}_{i}+D_{j})\}_{i=1}^{M}$ for each item $j$. Since sampling from $f(\bm{\theta}|Y_{1:t})$ only needs to be done once, computing the matrix $\Psi_{t}$ online is computationally efficient.
This practice of augmenting raw state variable with additional contextual features of the states ($\Psi_{T}$) echos the common strategy in the RL literature [Mnih2015, pmlr-v37-schaul15]: learning tends to be more stable and efficient when the state representation captures not only the current ability estimate but also characteristics of the item bank. In our setting, $\Psi_{t}$ offers a richer description of the predictive distributions across items given the current estimates, effectively serving as additional covariates for item selection. Empirically, we find that incorporating $\Psi_{t}$ substantially accelerates convergence toward the optimal item selection policy. Including $\Psi_{t}$ in the state representation is harmless, as it is a deterministic function of the posterior distribution; identical posteriors will always produce identical $\Psi_{t}$, ensuring that the policy remains consistent across equivalent states.
In summary, our proposed network structure can be summarized in Figure 2. To select the $(t+1)$-th item, our proposed neural net takes two inputs: the posterior parameters $\tilde{\bm{\xi}}_{t}$ and the prediction matrix $\Psi_{t}$, and outputs the selected item. Let $\phi_{2}:\mathbb{R}^{J\times Q}\to\mathbb{R}^{L_{2}}$ represent the encoder component that maps the matrix $\Psi_{t}\in\mathbbm{R}^{J\times Q}$ to $L_{2}$-dimensional space. Write $L=L_{1}+L_{2}$ and the concatenated outputs of $g_{1}(.)$ and $\phi_{2}(.)$ as $g_{1}(\tilde{\bm{\xi}}_{t})^{\prime},\phi_{2}(\Psi_{t})^{\prime}]^{\prime}\in\mathbb{R}^{L}$ , we then define the classification component of neural network as $\rho(.):\mathbb{R}^{L}\rightarrow\mathbb{R}^{J}$, which maps the concatenated outputs of $g_{1}(.)$ and $\phi_{2}(.)$ to a $J$-dimensional logit vector. Denote the final policy network as $\pi_{\phi}$ and recall that $\rho(.)$ represents the classification layer, our proposed network selects the item at time $t$ corresponding to the maximum value of the following function:
$\pi_{\phi}(\tilde{\bm{\xi}}_{t},\Psi_{t}):=\rho\circ\left\{[g_{1}(\tilde{\bm{\xi}}_{t})^{\prime},\phi_{2}(\Psi_{t})^{\prime}]^{\prime}\right\}.$
We implement both primary and target Q-networks as simple feed-forward multilayer perceptron and find that their performance is largely insensitive to the exact choice of $L_{1}$ and $L_{2}$, and does not require very large depth. All architectural details such as layer sizes, activations, and optimizer settings are provided in Appendix D. This robustness suggests that our empirical improvements stem from the CAT‐specific state design and augmentation rather than network complexity.
[…]
Figure 2: High level network. The shared encoder $\phi_{1}$ maps each tuple of posterior parameters to $\mathbb{R}^{L_{1}}$ and the sum yields the permutation invariant representation $g_{1}(\tilde{\bm{\xi}}_{t})$. The matrix $\Psi_{t}$ is encoded by $\phi_{2}$. The concatenated vector in $\mathbb{R}^{L}$ is passed to the classifier $\rho$ to select the $j$-th item (largest value in the $J$ logits).
5.2 Double Deep Q-Learning for CAT
The standard Q-learning algorithm often overestimates Q-values because the Q-network selects actions based on maximization operations, which are then used in network evaluation during training. To address this issue, we adopt the Double Q-learning approach [van_Hasselt_Guez_Silver_2016], which uses a primary network to select the best action and a separate target network to evaluate it. This separation results in more accurate Q-value approximations and leads to a more stable training process. Specifically, compared with standard Q-learning, Double Q-learning minimizes the same TD loss in (11) but replaces the target in (12) with
$y(s,a,r,s^{\prime})\;=\;r+\gamma\,Q_{\bar{w}}\!\big(s^{\prime},\,\operatorname*{arg\,max}_{a^{\prime}}Q_{w}(s^{\prime},a^{\prime})\big).$
Our proposed algorithm is presented in Algorithm table 1. Importantly, the offline training phase in Algorithm 1 relies only on the item parameters $(B,D)$ and simulated examinees drawn from a specified latent factor distribution (e.g., $\theta_{i}\sim\mathcal{N}(0,\mathbb{I}_{k})$), rather than on observed item response data from the item bank. If prior information about the distribution of future online examinees is available, the simulation distribution can be adapted accordingly. In this manuscript, we consistently use a standard multivariate normal distribution to ensure comparability across methods and experiments.
Note that for the Q-learning to converge to an optimal policy, it is essential to adopt an $\epsilon$-greedy policy, where the algorithm makes random item selections with probability $\epsilon$ and gradually decreases $\epsilon$ over the course of training. For simplicity, the state variable $s_{t}$ in Algorithm 1 is a compact representation of $(\tilde{\bm{\xi}}_{t},\Psi_{t})$ defined in Section 5.1. In practice, we can terminate the training when both the rewards and the validation loss stabilize, indicating that the neural network has well approximated the optimal policy.
Input: Item Parameters $(B,D)$, total Episodes $E$, $\epsilon\in[\underline{\epsilon},\bar{\epsilon}]$ with $\epsilon^{(0)}=\bar{\epsilon}$, buffer $\mathcal{H}$, target network update frequency $T$, batch size $M$.
Output: Learned Q-network $Q^{A}$ for item selection.
Initialize primary Q-network $Q^{A}$ and target Q-network $Q^{B}$, with weights $w^{A}_{0}=w^{B}_{0}$;
for $i=1$ to $E$ do
   Draw $\bm{\theta}_{i}$ from $\mathcal{N}(0,I_{k})$;
   Set initial state $s_{0}:=[\mathcal{N}(0,I_{k}),\Psi_{0}]$ and available item set $R_{0}=\{1,\cdots,J\}$;
 for $t=0$ to $H$ do
      Select action $a_{t}$ using $\epsilon$-greedy policy:
$a_{t}=\begin{cases}\text{Random Selection}&\text{with probability }\epsilon^{(t)}\\ \arg\max_{a\in R_{t}}\hat{Q}^{A}(s_{t},a)&\text{otherwise.}\end{cases}$
      Simulating response $y_{t}\sim\mathrm{Bernoulli}(\Phi(B_{t}^{\top}\bm{\theta}_{i}+d_{t}))$;
      Update available item set $R_{t+1}=R_{t}\setminus a_{t}$;
      Obtain new state $s_{t+1}:=[f(\theta|\bm{y}_{1:t+1}),\Psi_{t+1}]$
      compute reward $r_{t}$;
    if $r_{t}=0$ then
         Break and proceed to the next episode;
      end if
     Store transition $H_{i,t}:=(s_{t},a_{t},r_{t},R_{t},s_{t+1})$ into buffer $\mathcal{H}$;
    if $|\mathcal{H}|>\text{batch size }M$ then
         Randomly sample $M$ transitions $\{H_{i,t}\}_{j=1}^{M}$ from $\mathcal{H}$;
         Compute $a_{j}=\arg\max_{a\in R_{j}}\hat{Q}^{A}(s_{j}^{\prime},a)$;
         Compute target:
$y^{(j)}=\begin{cases}-1&r_{j}=0\\ -1+\gamma\hat{Q}^{B}(s_{j}^{\prime},a_{j})&\text{otherwise.}\end{cases}$
Perform batch gradient descent on loss: $\sum_{j=1}^{M}(y^{(j)}-\hat{Q}^{A}(s_{j},a_{j}))^{2}$;
      end if
     Decrease $\epsilon^{(t)}$ if $\epsilon^{(t)}\geq\underline{\epsilon}$;
   end for
  Update target network $Q^{B}$ such that $w_{t}^{A}=w_{t}^{B}$ if $i$ is a multiple of $T$;
end for
Algorithm 1 Double Q-Learning Algorithm for Bayesian CAT (offline)
5.3 Remark on Our Choice of Reward Function
Recall that our choice of reward function is defined in Equation (7). The simple 0-11 reward structure offers several advantages. It is interpretable as it directly minimizes the number of items required to terminate assessment tasks and prioritizes the main factors of interest. Additionally, it has demonstrated success in related applications, such as in designing deep adaptive learning system [doi:10.3102/10769986221129847]. Since the cumulative reward is constrained to integer values between H-H and 1-1, the action-value function can be more reliably approximated by the neural network.
Our double Q-learning algorithm enjoys theoretical convergence guarantees to the optimal policy [van_Hasselt_Guez_Silver_2016] and can accommodate arbitrary reward specifications. It is known that myopic policy cannot outperform the principled optimal policy obtained through RL. However, under certain reward structures, a greedy one-step lookahead (myopic) policy can perform surprisingly well: [chan2009stochastic] show that for a class of stochastic depletion problems, the myopic policy can attain at least $50\%$ of the optimal expected reward. In Appendix E, we establish an analogous result for CAT but also show that the 011 reward used in our framework violates the key assumptions required for such guarantees. Because our reward is explicitly designed to minimize test length, this violation underscores the value of the 011 reward function for adaptive testing.
6 Simulation
6.1 Simulation Design
To evaluate our approach under a challenging and realistic simulation setting, we randomly generated a $5$-factor, $150$-item factor loading matrix $B\in\mathbbm{R}^{150\times 5}$. Each column of $B$ was initialized by randomly permuting $150$ equally spaced values, with the magnitude of the discrimination parameters constrained to lie within $[0.3,3]$. We then imposed a lower triangular structure to ensure identifiability. To better reflect practical datasets, where items rarely load on all five factors, each item was required to load on the first factor and on at most two additional factors beyond the first. A visualization of the loading matrix can be found in Appendix F. Item intercepts were independently drawn from $\text{Unif}(-1.5,1.5)$.
The goal of the simulation is to accurately estimate the first three factors ($\mathcal{K}=\{1,2,3\}$), while accounting for the presence of factors $4$ and $5$. Leveraging our proposed direct sampling approach outlined in Section 4, we implemented our double Q-learning algorithm alongside all existing methods discussed in Section 3.2. These include the EAP approach (Equation 2), the Max Pos approach (Equation 3), the MI approach (Equation 5), and the Max Var approach (Equation 6). We further modified all baseline information-based criteria so that they also target only the prioritized subset of factors $\mathcal{K}=\{1,2,3\}$. Specifically, the EAP, Max Pos, and MI rules are now written as integrals over $\theta_{\mathcal{K}}$ rather than all $5$ factors. These adjustments ensure that all competing CAT algorithms are tuned to the same estimation target, making the simulation comparisons fair and directly comparable.
For performance evaluation, we generated $N=500$ online examinees with latent traits $\theta_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,I_{5})$. For each examinee, we administered a $50$-item test using each of the CAT algorithms. We then compared performance in terms of posterior variance reduction, mean-squared error (MSE), termination efficiency, and item exposure rates based on these $500$ adaptive testing sessions.
For the Q-learning algorithm, we trained the Q-network for $80,000$ episodes, with exploration parameter $\epsilon$ decreasing linearly from $0.99$ to $0.01$ over $700,000$ steps. We set a sufficient large $H=60$ items, with discount factor $\gamma=0.95$. To make sure the double Q-learning algorithm indeed converges to the optimal policy $\pi^{*}$, we provide further details on the training dynamics of our deep Q-learning algorithm, illustrating the increase in rewards until convergence in Appendix F.2.
6.2 Simulation Results
Figure 3: Number of Items Versus Cumulative Percentage of Completed Tests
To evaluate termination efficiency, we start with a standard multivariate Gaussian prior with unit marginal variances and terminate the test once the maximum posterior variance across all three target factors in $\mathcal{K}$ falls from 1 to below $\tau^{2}<0.16$. In practice, practitioners may select the threshold $\tau^{2}$ based on the specific requirements of their application. Based on $500$ simulated adaptive testing sessions, Figure 3 illustrates how the percentage of completed test sessions increases as more items are administered. Faster growth rate of the completion percentages indicates faster posterior variance reduction. Notably, the purple Q-learning curve consistently grows above all curves, demonstrating significant testing efficiency gains. As summarized in the second column of Table 1, Q-learning requires an average of only $21.5$ items to reach the termination criterion for all $3$ targeted factors, outperforming all other methods.
Table 1: Comparison of Win shares (W.S), Termination, and Computation
Algorithm Avg Termination (items) W.S dim1 W.S dim2 W.S dim3 Avg Time (s/item)
EAP 31.7 17.0% 16.4% 13.6% 0.037
Max Pos 34.0 16.8% 16.2% 18.6% 0.055
MI 23.2 19.0% 21.4% 20.2% 0.082
Max Var 25.7 23.6% 22.6% 23.6% 0.038
Q-learning 21.5 23.6% 23.4% 24.0% 0.064
To assess early-stage performance, we compute the MSEs after $20$ items for each test sessions and summarize results using win shares, defined as the percentage of test takers for whom each method achieves the lowest MSE after $20$ items. These win shares are reported across all three latent dimensions in Table 1. Notably, Q-learning achieves the best win shares across all three factors. This advantage at early stages is consistent with the training objective of Q-learning, which explicitly rewards rapid posterior variance reduction and early test termination, with the average stopping time occurring around $20$ items.
Table 2 reports the evolution of MSEs as a function of test length up to $50$ administered items. Consistent with the win share analysis, Q-learning exhibits competitive performance during the early stages of testing. In particular, at $T=20$ and $T=30$, Q-learning attains the smallest MSEs for Factors $2$ and $3$. As the test length increases, performance differences across methods diminish, and Q-learning achieves MSEs comparable to those of MI and MaxVar. This pattern is expected, as the Q-learning policy is trained primarily on short-horizon testing trajectories: since most simulated tests terminate before $30$ items, the algorithm is exposed less frequently to long-horizon simulations during training. Additional visualizations of the MSE trajectories are provided in Appendix F.3.
Table 2: MSEs between Posterior Mean and Ground Truth for the first three latent factors as a function of test length.
Number of items
Factor Algorithm 10 20 30 40 50
1 EAP 0.175 0.123 0.090 0.073 0.062
MaxPos 0.194 0.128 0.104 0.083 0.062
MI 0.180 0.100 0.066 0.055 0.050
MaxVar 0.186 0.080 0.073 0.059 0.056
Q-Learning 0.221 0.096 0.071 0.060 0.054
2 EAP 0.362 0.214 0.164 0.125 0.105
MaxPos 0.354 0.221 0.166 0.129 0.103
MI 0.223 0.150 0.120 0.102 0.086
MaxVar 0.214 0.141 0.118 0.092 0.087
Q-Learning 0.282 0.138 0.111 0.096 0.087
3 EAP 0.317 0.226 0.152 0.112 0.085
MaxPos 0.324 0.207 0.139 0.115 0.092
MI 0.213 0.127 0.103 0.096 0.089
MaxVar 0.234 0.136 0.110 0.099 0.096
Q-Learning 0.233 0.114 0.098 0.091 0.085
Figure 4: Distributions of Item Exposure Rates
Figure 4 summarizes the distribution of item exposure rates across the $150$ item bank for each CAT algorithm, where the exposure rate of an item is defined as the proportion of examinees whose test includes that item at least once. With test length $H=50$ and bank size $J=150$, the expected average exposure is $H/J\approx 0.33$, and all five algorithms yield mean and median exposure rates close to this benchmark. Differences across methods primarily arise in the tails of the exposure distribution. In particular, the EAP and Q-learning policies exhibit slightly higher upper-quartile and maximum exposure rates, indicating a more selective use of highly informative items. Importantly, no method concentrates exposure excessively on a small subset of items, suggesting that all algorithms maintain reasonable item utilization in this simulation setting. For the Q-learning approach, incorporating explicit exposure control mechanisms into the learning objective, such as penalizing repeated use of highly exposed items, is a natural extension but is beyond the scope of the present study.
The last column of Table 1 reports the average online item selection time, highlighting the significant computational advantages of our direct sampling approach. Even the most computationally intensive MI approach requires only $0.082$ seconds per item selection, via our proposed posterior reweighting strategy. Q-learning adds only a single feed-forward pass per selection, keeping online latency under a few milliseconds on standard hardware without GPU acceleration. While training the Q-network offline using Algorithm 1 took approximately $30$ hours for this exercise, this is a one-time investment that can be accelerated if GPUs are available; thereafter, the virtually instantaneous online policy makes the approach highly practical for real‐time CAT deployment.
7 Cognitive Function Measurements
7.1 pCAT-COG Data and Experiment Design
We revisit the problem of designing a deep CAT system for the pCAT-COG study, as outlined in Section 2. Since item response data for all $N=730$ examinees across $J=57$ items are available, we can directly use real item responses during evaluation rather than simulating testing sessions. In this experiment, all $N=730$ examinees are treated as online examinees. All CAT algorithms, including Q-learning, have access only to the estimated item parameters from the pCAT-COG study and do not observe the examinees’ binary item responses beyond those revealed sequentially during adaptive administration.
Given that pCAT-COG is designed to measure global cognitive ability (first column in Figure 1) while accounting for five cognitive subdomains, we specified the Q-learning reward function to reduce posterior variance for the primary factor ($\mathcal{K}=\{1\}$). This demonstrates the flexibility of our CAT system, as it can be tailored to the specific cognitive assessment needs. As in Section 6, we also modified all baseline CAT algorithms so that they also target only the primary factor by only integrating over the primary dimension for fair comparison.
Finally, this real-data experiment also serves as a robustness check for the proposed Q-learning approach. As described in Algorithm 1, the policy is trained using simulated examinees with latent factors drawn from a standard multivariate normal distribution. In contrast, the empirical factor correlations in pCAT-COG need not follow this distributional assumption. Evaluating Q-learning on real response data therefore provides evidence that the learned policy remains effective when the latent factor structure deviates from the training distribution used in simulation.
7.2 Results
The left subplot of Figure 5 shows that Q-learning again achieves the fastest test termination compared to other methods. As before, the test is dynamically terminated when the posterior variance drops from $1$ to below $\tau^{2}=0.16$, and the cumulative percentage of completed test sessions is computed over all $730$ examinees. The second column of Table 3 shows that Q-learning reaches the desired posterior variance reduction threshold after only an average of $11.2$ items.
For further comparison, we consider adaptively selecting $20$ items for each test taker without dynamic termination. We chose the number $20$ because the Max Pos approach required the largest average of $19.8$ items for termination. The third column of Table 3 shows that the Q-learning approach achieves the highest $0.959$ correlation between the estimated primary factor posterior means and the ground truth after only $20$ items. Additionally, Q-learning approach attains the highest win shares in estimating the primary dimension across all $N=730$ examinees after $20$ items. As defined in Section 6, Win shares denote the proportion of examinees for whom a CAT method attains the smallest mean-squared error. Since all $57$ items are ultimately administered in the real-data study, exposure rates are degenerate under full-length testing. We therefore report item exposure patterns based on the first $30$ administered items in Appendix G.1.
Additionally, the right subplot of Figure 5 highlights the rapid decay of mean squared error (MSE) in estimating the primary dimension with Q-learning. As summarized in Table 4, Q-learning achieves the smallest MSE at the early stages when $T=10$ and $T=20$. At longer horizons when $T=30$ and $T=40$, the gains in MSE for the Q-learning approach diminishes. This behavior is consistent with the simulation findings in Section 6 and reflects the reward specification used for Q-learning, which emphasizes rapid early-stage variance reduction. Because most simulated training trajectories generated in Algorithm 1 terminate well before 20 items, such performance behavior for Q-learning is expected.
(a) Test Termination Speed
(b) Test MSE Accuracy
Figure 5: pCAT-COG: Primary Factor Posterior Variance Reduction (Left) and Estimation Accuracy (Right)
Table 3: Comparison of termination efficiency, primary-factor accuracy, and computation for pCAT-COG.
Algorithm Avg Termination (items) Correlation Win shares Avg Time (s/item)
EAP 15.1 0.956 17.7% 0.020
Max Pos 18.5 0.933 16.7% 0.030
MI 12.6 0.959 21.8% 0.038
Max Var 13.0 0.951 20.9% 0.021
Q-Learning 11.2 0.959 22.9% 0.025
  • Note. Primary factor correlation and win shares (W.S) are computed after 20 administered items. Win shares denote the proportion of examinees for whom a method attains the smallest mean-squared error for the primary factor at this stage.
Table 4: Mean-squared errors (MSEs) as a function of test length
Algorithm 10 items 20 items 30 items 40 items
EAP 0.183 0.079 0.029 0.010
Max Pos 0.217 0.113 0.067 0.016
MI 0.165 0.070 0.032 0.014
Max Var 0.165 0.085 0.038 0.014
Q-Learning 0.123 0.072 0.038 0.016
This experiment highlights the effectiveness of our Q-learning approach in high-dimensional cognitive function measurement. Unlike approaches that learn a fixed test design offline [Krantsevich2023], our framework enables fully dynamic item selection, maximizing the utilization of the item bank by exploring diverse testing trajectories. This termination efficiency is particularly valuable in cognitive assessment, where item development is costly, and can help mitigate practice effects and preserves items for future use.
8 Discussion
This work advocates for a deep reinforcement learning perspective in the design of multidimensional computerized adaptive testing systems. We make two key contributions to the existing CAT literature: (1) a computational advancement that accelerates existing online item selection rules within a flexible Bayesian MIRT framework, and (2) a novel reinforcement learning-based approach that mitigates myopic decision-making and prioritizes the assessment of primary factors of interest.
Theorem 4.2 not only provides an efficient parameterization of the latent factor posterior distributions for our proposed deep Q-learning approach, but also improves existing CAT item selection algorithms, as detailed in Section 4. Leveraging direct sampling from unified skew-normal distributions, our methodology scales efficiently with a large number of factors and items, achieving near-instantaneous online selection by circumventing MCMC sampling and data augmentation. Additionally, our approach naturally extends to fully Bayesian item selection by accounting for uncertainties in item parameters, which is an essential consideration when the item bank is not well-calibrated as detailed in Appendix C. Throughout the testing trajectory, our approach precisely characterizes the evolution of posterior distributions at each time step, providing a more robust measurement process beyond point estimates with difficult-to-compute standard errors.
Another key contribution of this work is the development of a robust deep double Q-learning algorithm with a customized reward structure that directly minimizes test length. As demonstrated in both simulations and real-data studies, our Q-learning algorithm consistently achieves the fastest posterior variance reduction while rapidly decreasing estimation bias. Moreover, its flexible reward function allows adaptation to different testing objectives, providing a principled framework for designing customized tests and overcoming the myopic nature of traditional CAT methods.
Our work also offers practical guidance for selecting the appropriate item selection algorithms. Unlike the existing heuristic rules, one limitation of our deep CAT system is the requirement of offline training (Algorithm 1) before online deployment. For the pCAT-COG study, offline training took approximately $12$ hours on a single GPU as an one-time investment, but no GPU is needed for subsequent online item selection. Even when offline training is undesirable, our framework significantly accelerates existing methods. Experiments show that MI (Equation 5) and our maximizing variance (Equation 6) approaches often outperform other online item selection rules, echoing the findings presented in [selection_survey].
A promising direction for future work is to address several limitations of the current Q-learning framework for CAT. First, the learned policy is tied to a specific item bank and must be retrained when the items are substantially modified or replenished, which may limit its immediate applicability in settings with frequent item updates. Second, as a model-free approach, Q-learning can be computationally expensive to train, particularly for long-horizon testing scenarios in which optimal policies depend on extended future trajectories. Finally, while the proposed method yields reasonable item exposure patterns in our experiments, additional methodological development is needed to incorporate explicit exposure control mechanisms directly into the learning objective. Addressing these challenges would further enhance the practicality and robustness of RL–based CAT systems.
Another exciting avenue is to explore alternative reward structures for our Q-learning algorithm. While the 0-1 reward structure is interpretable and stabilizes Q-network training, it provides sparse feedback, which may limit empirical performance. Alternative designs could incorporate intermediate rewards or increased penalization as more items are administered. An interesting theoretical direction is to characterize when reinforcement learning approaches are particularly advantageous compared to myopic methods. This requires careful assumptions about the reward function and the item bank’s properties, providing deeper insights into the trade-offs between RL and traditional item selection strategies.
Supplementary Material: Deep Computerized Adaptive Testing
Appendix A Mutual Information as Prediction Uncertainties
We observe all the heuristic Bayesian item selection methods discussed in this manuscript account for prediction uncertainty. In particular, the mutual information item selection rule can be rewritten as follows:
$\arg\max_{j_{t}\in R_{t}}I_{M}(\theta,y_{j_{t}})$ $=\arg\max_{j_{t}\in R_{t}}\sum_{y_{j_{t}}=0}^{1}\int_{\theta}f(\theta,y_{j_{t}}|Y_{t-1})\log\frac{f(\theta,y_{j_{t}}|Y_{t-1})}{f(\theta|Y_{t-1})f(y_{j_{t}}|Y_{t-1})}d\theta$
$=\arg\max_{j_{t}\in R_{t}}\sum_{y_{j_{t}}=0}^{1}\int_{\theta}f(y_{j_{t}}|\theta)f(\theta|Y_{t-1})\log\frac{f(\theta,y_{j_{t}}|Y_{t-1})}{f(\theta|Y_{t-1})f(y_{j_{t}}|Y_{t-1})}d\theta$
$=\arg\max_{j_{t}\in R_{t}}\sum_{y_{j_{t}}=0}^{1}\int_{\theta}f(y_{j_{t}}|\theta)f(\theta|Y_{t-1})\log\frac{f(y_{j_{t}}|\theta)}{f(y_{j_{t}}|Y_{t-1})}d\theta$
$=\arg\max_{j_{t}\in R_{t}}\int_{\theta}[\Phi(B_{j_{t}}^{\prime}\theta+D_{j_{t}})\log\frac{\Phi(B_{j_{t}}^{\prime}\theta+D_{j_{t}})}{c_{j_{t}}}+$ (18)
$\hskip 60.00009pt(1-\Phi(B_{j_{t}}^{\prime}\theta+D_{j_{t}}))\log\frac{(1-\Phi(B_{j_{t}}^{\prime}\theta+D_{j_{t}}))}{(1-c_{j_{t}})}]f(\theta|Y_{t-1})d\theta.$
Equation (A) suggests the mutual information criterion selects the item $j_{t}$ that has the largest expected KL divergence between the Bernoulli distributions parametrized by $\Phi(B_{jt}^{\prime}\theta+d_{jt})$ and $c_{jt}$, weighted by the current posterior $f(\theta|Y_{t-1})$.
Observe that mutual information essentially favors the item $j_{t}$ that has the largest prediction uncertainties around its prediction mean $c_{j_{t}}$, quantified by the KL divergence. To compare equation (A) with equation (6), note we have replace the integral of KL divergence from $c_{j_{t}}$ with variances $(\Phi(B_{j_{t}}^{\prime}\theta+D_{j_{t}})-c_{j_{t}})^{2}$ in (6). This suggests that our proposed approach shares similar theoretical properties with the mutual information criterion, but offers a much simpler formula for online computation, avoiding posterior reweighting and log transformations. Empirically, our proposed item selection rule achieves comparable performance to mutual information but with significantly reduced computational time.
Appendix B Sequential Sampling in CAT
B.1 Proof of Theorem 4.2
For any $T\geq 1$, we may write the liblackkelihood of item response after administering $T$ items under the two-parameter probit MIRT model as follows:
$\prod_{t=1}^{T}\Phi(B_{j_{t}}^{\prime}\bm{\theta}+D_{j_{t}})^{y_{j_{t}}}(1-\Phi(B_{j_{t}}^{\prime}\bm{\theta}+D_{j_{t}}))^{1-y_{j_{t}}}$ $=\prod_{t=1}^{T}\Phi\{(2y_{j_{t}}-1)(B_{j_{t}}^{\prime}\bm{\theta}+D_{j_{t}})\}$
$=\Phi_{T}\{((C_{1}\bm{\theta}+C_{2});\mathbbm{I}_{T}\}.$
Given the standard multivariate normal prior $\theta\sim N(0,\mathbbm{I}_{K})$, the posterior can be expressed as follows:
$\pi(\bm{\theta}|Y_{1:T},B_{1:T},D_{1:T})$ $\propto\phi_{K}(\bm{\theta};\bm{0},\mathbbm{I}_{k})\Phi_{T}\{((C_{1}\bm{\theta}+C_{2});\mathbbm{I}_{T}\}$
$=\phi_{K}(\bm{\theta};\bm{0},\mathbbm{I}_{k})\Phi_{T}\{C_{3}^{-1}((C_{1}\bm{\theta}+C_{2});C_{3}^{-1}C_{3}^{-1}\}$
$=\phi_{K}(\bm{\theta};\bm{0},\mathbbm{I}_{k})\Phi_{T}\{C_{3}^{-1}C_{2}+C_{3}^{-1}C_{1}\bm{\theta};C_{3}^{-1}C_{3}^{-1}\}$
To identify the exact posterior parameters, we draw back to our attention to the probability kernel of the unified skewed-normal distribution as defined in Definition 4.1. We observe that the posterior density above aligns with the SUN density. Specifically, if $\bm{\theta}\sim\operatorname{SUN}_{K,T}(\mu,\Omega,\Delta,\gamma,\Gamma)$, its density is
$\phi_{K}\!\big(\bm{\theta};\mu,\Omega\big)\,\frac{\Phi_{T}\!\left(\gamma+\Delta^{\top}\bar{\Omega}^{-1}\omega^{-1}(\bm{\theta}-\mu)\,;\,\Gamma-\Delta^{\top}\bar{\Omega}^{-1}\Delta\right)}{\Phi_{T}(\gamma;\Gamma)}.$
We can interpret this as a product of a $K$-variate Gaussian pdf in $\bm{\theta}$ and a $T$-variate Gaussian cdf in $\bm{\theta}$, with normalizing constant $\Phi_{T}(\gamma;\Gamma)$, since $\Phi_{T}(\gamma;\Gamma)$ is not a function of $\bm{\theta}$ . But this exactly matches our derived latent-factor posterior $\pi(\bm{\theta}\mid Y_{1:T},B_{1:T},D_{1:T})$ under the MIRT model. It remains to identify $(\mu,\Omega,\Delta,\gamma,\Gamma)$ by matching $\phi_{K}\!\big(\bm{\theta};\mu,\Omega\big)\Phi_{T}\!\left(\gamma+\Delta^{\top}\bar{\Omega}^{-1}\omega^{-1}(\bm{\theta}-\mu)\,;\,\Gamma-\Delta^{\top}\bar{\Omega}^{-1}\Delta\right)$ to
$\phi_{K}\!\big(\bm{\theta};\mathbf{0},\mathbf{I}_{K}\big)\,\Phi_{T}\!\left(C_{3}^{-1}C_{2}+C_{3}^{-1}C_{1}\bm{\theta}\,;\,C_{3}^{-1}(C_{3}^{-1})^{\top}\right).$
Matching the pdf part immediately yields $\mu_{\text{post}}=0_{K}$ and $\Omega_{\text{post}}=\mathbbm{I}_{K}$. To match the cdf part, we need to match $\gamma+\Delta^{\top}\bar{\Omega}^{-1}\omega^{-1}(\bm{\theta}-\mu)$ to $C_{3}^{-1}C_{2}+C_{3}^{-1}C_{1}\bm{\theta}$, which implies $\gamma_{\text{post}}=C_{3}^{-1}C_{2}$. It remains to solve the following two linear equation:
$\Delta^{\prime}\bar{\Omega}^{-1}\omega^{-1}=C_{3}^{-1}C_{1}$
$\Gamma-\Delta^{\prime}\bar{\Omega}^{-1}\Delta=C_{3}^{-1}C_{3}^{-1}.$
Recall that in definition 4.1, we let $\Omega=\omega\bar{\Omega}\omega$ be the decomposition of covariance matrix into correlation matrix. Since $\Omega_{\text{post}}=\mathbbm{I}_{K}$, we have $\bar{\Omega}=\omega=\mathbbm{I}_{K}$. It follows that solving the first equation yields $\Delta_{\text{post}}=C_{1}^{\prime}C_{3}^{-1}$. Plugging our solution of $\Delta_{\text{post}}$ in the second solution yields
$\Gamma_{\text{post}}=C_{3}^{-1}C_{3}^{-1}+C_{3}^{-1}C_{1}C_{1}^{\prime}C_{3}^{-1}=C_{3}^{-1}(C_{1}C_{1}^{\prime}+\mathbbm{I}_{T})C_{3}^{-1}.$
Finally, to show this is indeed a unified skewed-normal distribution, we need to show the matrix $\Omega^{*}$ as defined in Definition 4.1 is indeed a full-rank correlation matrix. To see this, we may decompose $\Omega^{*}$ as follows:
$\begin{bmatrix}C_{3}^{-1}(C_{1}C_{1}^{\prime}+\mathbbm{I}_{T})C_{3}^{-1}&C_{3}^{-1}C_{1}\\ C_{1}^{\prime}C_{3}^{-1}&\mathbbm{I}_{K}\\ \end{bmatrix}=\begin{bmatrix}C_{3}^{-1}&0\\ 0&\mathbbm{I}_{k}\\ \end{bmatrix}\times\begin{bmatrix}C_{1}C_{1}^{\prime}+\mathbbm{I}_{T}&C_{1}\\ C_{1}^{\prime}&\mathbbm{I}_{k}\\ \end{bmatrix}\times\begin{bmatrix}C_{3}^{-1}&0\\ 0&\mathbbm{I}_{k}\\ \end{bmatrix}$.
Observe that this is a decomposition of a correlation matrix, where the middle matrix above is the covariance matrix of a $(T+K)$ dimensional random vector $z_{1}^{\prime},z_{2}^{\prime}]^{\prime}$, where $z_{1}=C_{1}z_{2}+\epsilon$, $z_{2}$ is a $K$-dimensional standard multivariate Gaussian vector with identity covariance matrix, and $\epsilon$ is a $T$-dimensional standard multivariate Gaussian vector, independent of $z_{2}$.
B.2 Sequential Sampling Algorithms
For any Bayesian item selection algorithms or our proposed deep CAT framework, it is essential to sample from the latent factor posterior distributions $f(\theta|Y_{1:t})$ for any $t\geq 1$. Theorem 4.2 provides a direct sampling approach to perform such sequential sampling, hence circumventing the need for MCMC algorithms, which cannot be parallelized and require additional tuning and mixing time.
The key observation is that any unified skew-normal distribution $\theta\sim\operatorname{SUN}_{K,T}(\mu,\Omega,\Delta,\gamma,\Gamma)$ has the stochastic representation as $\theta\overset{d}{=}\mu+\omega(V_{0}+\Delta\Gamma^{-1}V_{1,-\gamma}),$ where $V_{0}\sim N(0,\bar{\Omega}-\Delta\Gamma^{-1}\Delta^{\prime})\in\mathbbm{R}^{K}$ and $V_{1,-\gamma}$ is obtained by component-wise truncation below $\gamma$ of a variate $N(0,\Gamma)\in\mathbbm{R}^{T}$. By plugging in the posterior parameters derived in Theorem 4.2, we have the representation
$\theta\sim V_{0}+C_{1}^{\prime}(C_{1}C_{1}^{\prime}+\mathbbm{I}_{t})^{-1}C_{3}V_{1,-\gamma}.$
Hence we can conduct efficient sequential posterior sampling using the following three steps when $t>1$:
  • Step One: Sample from multivariate normal distribution
    $V_{0}\sim N(0,\mathbbm{I}_{K}-C_{1}^{\prime}(C_{1}C_{1}^{\prime}+\mathbbm{I}_{T})^{-1}C_{1}.$
  • Step two: Leveraging the minimax tilting method [Botev_2016], sample from the zero-mean $T$-variate truncated multivariate normal distribution $V_{1,-\gamma}$ with covariance matrix $C_{3}^{-1}(C_{1}C_{1}^{\prime}+\mathbbm{I}_{T})C_{3}^{-1}$ and truncation below -$C_{3}^{-1}C_{2}$.
  • Step three: perform linear computation $V_{0}+C_{1}^{\prime}(C_{1}C_{1}^{\prime}+\mathbbm{I}_{T})^{-1}C_{3}V_{1,-\gamma}.$
Note for $t=1$, the sampling steps above can be modified into their univariate analog. Since steps one and step two are independent, one can simply perform i.i.d sampling from both $V_{0}$ and $V_{1}$ in parallel to draw large samples of $\theta\sim f(\theta|Y_{1:t})$ in the probit MIRT model. The $t\times t$ matrix inverse $(C_{1}C_{1}^{\prime}+\mathbbm{I}_{t})$ is easy to compute and only needs to be computed once.
Appendix C Fully Bayesian Item Selection
C.1 Generalizing to Fully Bayesian Item Selections
Another key advantage of our direct sampling approach is its ability to enable rapid fully Bayesian item selection. Since Section 3, we have assumed that the item bank is well-calibrated, allowing the factor loading matrix and the intercept to be treated as fixed. According to fully_bayesian, this assumption can lead to overly optimistic estimation of the latent traits in practice, and is only reasonable when large calibration datasets are available. A fully Bayesian framework offers a principled way to incorporate this uncertainty by integrating over the joint posterior of item parameters and latent traits. However, performing this integration in online item selection has long been considered intractable, limiting its practical adoption. Notably, fully_bayesian introduced an efficient MCMC method, but only for unidimensional item response theory model.
Our direct sampling approach extends fully Bayesian item selection to multidimensional IRT while avoiding the computational bottlenecks of MCMC. To illustrate, we consider the mutual information criterion in (15). Fully Bayesian inference involves obtaining posterior samples of item parameters $\{\bm{\xi}^{(m)}\}_{m=1}^{M}$, where $\bm{\xi}^{(m)}:=(B^{(m)},D^{(m)})$. The mutual information criterion in Equation (15) requires computing the posterior predictive probability:
$f(y_{jt}|Y_{1:(t-1)})=\int_{\bm{\xi}}\int_{\bm{\theta}}f(y_{j_{t}}|\bm{\theta})f(\bm{\theta}|\bm{\xi},Y_{1:(t-1)})d\bm{\theta}d\bm{\xi}.$ (19)
Similarly, the KL divergence term must integrate out the nuisance item parameters $\bm{\xi}$:
$\int_{\bm{\xi}}\int_{\bm{\theta}}f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}},\bm{\xi})\log\frac{f(y_{i_{t}}|\bm{\theta})}{f(y_{j_{t}}|Y_{1:(t-1)})}d\bm{\theta}d\bm{\xi}.$ (20)
By leveraging Theorem 4.2, we can exactly identify the posterior distributions for any arbitrary configuration of $f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}},\bm{\xi}^{(m)})$, hence enabling direct sampling in parallel. It follows Equations (19) and (20) are computationally feasible via Monte Carlo integration.
In contrast, a standard MCMC approach requires constructing $M$ independent Markov chains with additional data augmentation techniques [f3928584-48a1-3ab1-b404-2ec9ffdf51bb, Polson01122013], each targeting $f(\bm{\theta}|Y_{1:(t-1)},y_{j_{t}},\bm{\xi}^{(m)})$, with sampling only possible after all chains have reached convergence. This introduces a significant computational burden, making MCMC impractical for real-time item selection.
We evaluate the effectiveness of our direct sampling approach in enabling fully Bayesian online item selection. Unlike traditional methods that treat item parameters as fixed, the fully Bayessian approach explicitly accounts for their uncertainties, which can be substantial when the item response dataset is small or poorly calibrated. To illustrate, we generate a binary item response dataset $Y$ with $N=500$ examinees and $J=150$ items under a 3-factor probit MIRT model. The true factor loading matrix $B$ has dimensions $J\times 3$, and the intercept vector $D$ has $150$ elements.
To integrate item parameter uncertainty into online item selection, we fit a Gibbs sampler to the $N\times J$ item response data. We generated $5,000$ MCMC draws, retaining the last $500$ posterior samples of the item parameters $\Xi={\xi^{(m)}}_{m=1}^{500}$ after burn-in. As described in Section C.1, the fully Bayesian approach marginalizes over the joint distributions of the nuisance item parameters $\Xi$ and latent traits. We then implemented all item selection rules depicted in Sections 3.2 in a fully Bayesian manner.
Figure 6 illustrates the cumulative percentage of completed test sessions across all $500$ simulated examinees as more items are administered. The test was dynamically terminated when the posterior standard deviations across all three factors fell below $0.4$. Consistent with the experimental results in Section 6 of the manuscript, the mutual information method and our proposed Max Var method demonstrated superior posterior variance reduction. The second column of Table 5 shows that the mutual information method terminated after an average of $18.7$ items, whereas the EAP approach required $24.4$ items.
Figure 6: Fully Bayesian: Number of Items Versus Posterior Variance Reduction
Table 5: Comparison of Winshares (W.S), Termination, and Computation
Algorithm Avg Termination (items) W.S dim0 W.S dim1 W.S dim2 Avg Time (s/item)
EAP 24.4 22.2% 19.2% 21.8% 4.10
Max Pos 24.2 17.4% 18.0% 21.0% 4.10
MI 18.7 32.8% 31.0% 27.8% 4.11
Max Var 19.2 27.6% 31.8% 29.4% 4.10
Beyond variance reduction, the mutual information and maximizing prediction variance methods also exhibited superior estimation accuracy, as shown in Figure 7, which tracks the decline in mean squared error (MSE) between the posterior mean and the oracle posterior mean. Since estimation errors stabilized around $H=30$ items, we compared MSEs between the posterior mean at $H=20$ and the oracle posterior mean, and then summarized the win shares for each item selection rule across all $500$ examinees and all three dimensions in Table 5.
Figure 7: Fully Bayesian: MSE Between Posterior Mean and Oracle Posterior Mean
Finally, we emphasize that even in a fully Bayesian framework, where nuisance parameters are integrated rather than fixed, our approach remains computationally efficient. Across all item selection rules, online selection required an average of just $4.1$ seconds per test session on a personal laptop, highlighting the scalability of our direct sampling method for Bayesian MCAT applications. Readers may notice that $4.1$ seconds is significantly longer than the item selection time reported in the experimentation section, where item parameters are treated as fixed. This discrepancy arises because fully Bayesian MCAT requires sampling from $M=500$ distinct unified-skew normal distributions for each item selection, whereas the standard case involves sampling from a single fixed distribution. This also explains the minimal differences in item selection time across algorithms, as the primary computational bottleneck lies in sampling from these distributions. When more CPUs are available, we can certainly sampling from all $M$ distributions in parallel, and hence further improving the $4.1$ seconds benchmark.
Appendix D Neural Network Design Details
Both the primary and target Q–networks share the same modular feed-forward design. At each time $t$, the posterior parameters $\tilde{\xi}_{t}$ are split into two streams: a permutation-invariant path $\phi_{1}$ (three $256$-unit ReLU layers) that processes the main state input. In parallel, a RowwiseNetwork $\phi_{2}()$ network applies independently to each of the $J$ prediction-quartile feature vectors yields item-wise predictive values, which are then mapped to $256$ dimensional features. The two $256$-dimensional vectors are then merged by a small combiner MLP, which is then mapped to the final layers with $J$ nodes through the classification network $\rho$. All linear layers use Xavier initialization; activations are ReLU; the final output is masked to -$\infty$ for unavailable items.
Table 6: Layer‐by‐layer specification of OnlineQNetworkV (and its submodules).
Module Input Dim Output Dim Activation
$\phi_{1}$ (Phi1Network) $S$ 256 ReLU
256 256 ReLU
256 256 ReLU
$\phi_{2}$ (RowwiseNetwork) (for $\Psi_{t}$) 11 256 ReLU
256 256 ReLU
256 1 Linear $\times J$ rows
$\phi_{2}$ (SmallPhi2Network) $J$ 256 ReLU
256 256 ReLU
256 256 Linear
$\rho$ (Classification) 256 $J$ ReLU
$J$ $J$ Linear
Notes:
  • $S=$ total scalar‐state dimension; Each new item selection introduced $(k+2)$ dimensional tuple
  • The RowwiseNetwork applies its three layers independently to each of the $J$ rows of the $(J\times 11)$ $\Psi_{t}$ matrix. There are $11$ columns, because we include mean, variance, and the $10\%$ quantile increments from $10\%$ to $90\%$. Because we process each row independently, the output for the RowwiseNetwork is $J$ dimensional.
  • All weight matrices are initialized with Xavier uniform; no dropout is used.
  • Optimizer: Adam with learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$.
Appendix E Potential of the 0-11 Reward
In chan2009stochastic, it is shown that for stochastic depletion problems, a myopic policy achieves at least $50\%$ of the expected reward of the optimal policy. A similar result can be proven for CAT under specific reward structures. Although this can be interpreted as a somewhat negative result for the reinforcement learning approach, we emphasize our choice of the 0-11 reward does not conform to the stochastic depletion problem, and hence is not applicable to theorem below.
To illustrate this, we extend the state space $\mathcal{S}$ with a time variable, defining $\overline{\mathcal{S}}=\{(f(\bm{\theta}\mid Y_{1:t}),t):t\in[H]\}$. Let $\pi^{*}(\overline{s})$ represent the optimal policy, and $\pi^{m}(\overline{s})$ the myopic policy that maximizes the next-step expected reward. Define the mapping $\hat{G}_{y_{a}}:\overline{s}\rightarrow\overline{s^{\prime}}$ as the posterior update after observing the binary response random variable $y_{a}$, with an incremented time step, where $\overline{s}=(f(\bm{\theta}|Y_{1:t}),t)$ and $\overline{s^{\prime}}=(f(\bm{\theta}|Y_{1:t},y_{a}),t+1)$. Similarly, define $\tilde{G}_{y_{a}}:\overline{s}\rightarrow\overline{s^{\prime}}$ as the posterior update after observing $y_{a}$ while keeping the time step fixed, where $\overline{s}=(f(\bm{\theta}|Y_{1:t}),t)$ and $\overline{s^{\prime}}=(f(\bm{\theta}|Y_{1:t},y_{a}),t)$.
To demonstrate the near-optimality of the myopic policy, chan2009stochastic introduces two key assumptions: “value function monotonicity” and “immediate rewards”. These assumptions can be adapted to the CAT setting as follows:
  • Immediate Rewards: for any time stamp $t<H$, suppose the myopic policy $\pi^{m}$ chooses item $a$, then we must have
    $V_{\pi^{*}}(\overline{s})\leq E[R(\overline{s},a,\bar{s}^{\prime})]+V_{\pi^{*}}(\tilde{G}_{y_{a}}(\overline{s})).$ (21)
    Essentially, the property says given a free item (without increasing the time step) chosen by the myopic policy would not deteriorate the optimal rewards.
  • Value Function Monotonicity: for any time stamp $t<H$, suppose the optimal policy would choose item $a_{1}$ yet the myopic policy would chooses item $a_{2}$, then
    $V_{\pi^{*}}(\hat{G}_{y_{a_{2}}}(\tilde{G}_{y_{a_{1}}}(\overline{s})))\leq V_{\pi^{*}}(\hat{G}_{y_{a_{2}}}(\overline{s})).$ (22)
    Note both sides of inequality refer to the optimal value function evaluated at the posterior distribution at time step $t+1$.
In Theorem E.1, we establish that a reinforcement learning approach to CAT cannot significantly outperform the one-step-lookahead greedy policy when the reward function satisfies the two assumptions outlined above.
Theorem E.1.
Consider the finite horizon CAT problem with $H$ steps. Suppose the assumptions of immediate rewards in (21) and value function monotonicity in (22) hold, we have $V_{\pi^{*}}(\overline{s})\leq 2V_{\pi^{m}}(\overline{s})$ for all $\overline{s}\in\overline{\mathcal{S}}$.
Proof.
We proceed by induction. Since the myopic and the optimal policy would agree on time t=H1t=H-1, the inequality holds trivially for the base case. For the induction step, we assume the claimed inquality holds for all $t^{\prime}$ such that $H>t^{\prime}>t$. Consider the time horizon $t<t^{\prime}$ and let $a_{1}$ represent the item chosen by the optimal policy and $a_{2}$ the item chosen by the myopic policy. The inequality holds trivially for $a_{1}=a_{2}$. Hence, we consider the case when $a_{1}\neq a_{2}$ and let $y_{a_{1}}$ and $y_{a_{2}}$ represent the binary random variables for the item response. For given extended state $\overline{s}$, we use $\overline{s^{\prime}}$ to denote the next step after selecting an item. Then we have
$V_{\pi^{*}}(\overline{s}|y_{a_{1}},y_{a_{2}})$ $=E[R(\overline{s},a_{1},\overline{s^{\prime}})|y_{a_{1}}]+V_{\pi^{*}}^{*}(\hat{G}_{y_{a_{1}}}(\overline{s}))$
$\leq E[R(\overline{s},a_{1},\overline{s^{\prime}})|y_{a_{1}}]+E[R(\overline{s},a_{2},\overline{s^{\prime}})|y_{a_{2}}]+V_{\pi^{*}}(\tilde{G}_{y_{a_{2}}}(\hat{G}_{y_{a_{1}}}(\overline{s})))$ (23)
$=E[R(\overline{s},a_{1},\overline{s^{\prime}})|y_{a_{1}}]+E[R(\overline{s},a_{2},\overline{s^{\prime}})|y_{a_{2}}]+V_{\pi^{*}}(\hat{G}_{y_{a_{1}}}(\tilde{G}_{y_{a_{2}}}(\overline{s})))$
$\leq E[R(\overline{s},a_{1},\overline{s^{\prime}})|y_{a_{1}}]+E[R(\overline{s},a_{2},\overline{s^{\prime}})|y_{a_{2}}]+V_{\pi^{*}}(\hat{G}_{y_{a_{2}}}(\overline{s})))$ (24)
$\leq E[R(\overline{s},a_{1},\overline{s^{\prime}})|y_{a_{1}}]+E[R(\overline{s},a_{2},\overline{s^{\prime}})|y_{a_{2}}]+2V_{\pi^{m}}(\hat{G}_{y_{a_{2}}}(\overline{s}))).$ (25)
Note we applied the immediate rewards assumption in equation (23), the value function monotonicity assumption in equation (24), and the induction hypothesis in equation (25). Now, we may take expectation over $y_{a_{1}}$ and $y_{a_{2}}$ on both side of inequality and obtain:
$V_{\pi^{*}}(\overline{s})$ $\leq E[R(\overline{s},a_{1},\overline{s^{\prime}})]+E[R(\overline{s},a_{2},\overline{s^{\prime}})]+2V_{\pi^{m}}(\hat{G}_{y_{a_{2}}}(\overline{s})))$
$\leq 2(E[R(\overline{s},a_{2},\overline{s^{\prime}})]+V_{\pi^{m}}(\hat{G}_{y_{a_{2}}}(\overline{s}))))=2V_{\pi^{m}}(\overline{s}).$
The last inequality is due to the fact the $\pi^{m}$ is the myopic policy that maximizes the next step reward. ∎
Theorem E.1 can be seen as a limitation of the reinforcement learning approach when the reward structure aligns with the stochastic depletion problem, as it implies that a simple myopic policy provides a 2-approximation to the optimal policy. However, it also highlights the advantage of the simple 0-11 reward function as the value function monotonicity is unlikely to hold, rendering the theorem inapplicable. Specifically, the expected posterior variance of $\hat{G}_{y_{a_{2}}}(\tilde{G}_{y_{a_{1}}}(\overline{s})))$ is expected to be smaller than that of $\hat{G}_{y_{a_{2}}}(\overline{s})$. Consequently, fewer items are expected for the posterior $\hat{G}_{y_{a_{2}}}(\tilde{G}_{y_{a_{1}}}(\overline{s})))$ to reach the minimum variance threshold, which implies $V_{\pi^{*}}(\hat{G}_{y_{a_{2}}}(\tilde{G}_{y_{a_{1}}}(\overline{s})))\geq V_{\pi^{*}}(\hat{G}_{y_{a_{2}}}(\overline{s}))$.
Appendix F More Analysis on Simulation
F.1 Factor Loading Matrix Visualization
We visualize the factor loading matrix in our simulation exercise in Figure 8. Note all the discrimination parameters are bounded between (3,3)(-3,3) and the matrix is relatively sparse: all the items are loaded on the first factor, but are only allowed to load for at most two more factors. Although identification of the factor loading matrix is not a primary concern in CAT, as it is typically treated as fixed, we purposefully imposed a lower-triangular structure to ensure identifiability and create a more realistic simulation setting.
Figure 8: simulated-factor-loading
F.2 Double Q-Learning Training Dynamics
In Figure 9, we present the metric illustrating the training dynamics of our Q-learning algorithm. The left subplot displays the average episode rewards, computed over every 500 training episodes, showing that rewards began converging around episode 25,000 at approximately 23-23. The reward is bounded within [40,20][-40,-20]. Given our definition of the 010-1 reward, the implication is that a random policy would take $40$ items for the test to converge. After training, it would only take around $20$ items.
Since we save the primary neural network every 1000 episodes, we select the checkpoint corresponding to the highest average reward for offline deployment in future item selection tasks. In this case, the network at episode 72,000 is chosen, as it achieves the peak average reward of approximately 21-21. This model selection strategy is consistently applied across all experiments in Section 5.
Figure 9: Training Dynamics for the Simulation Experiment in Section 5.1
F.3 Matching the Oracle Posterior
Recall the ”oracle posterior distribution” of a test taker is defined as the posterior obtained if the given test taker had answered all items in the item bank. To obtain the oracle distributions for each of the $500$ test taker in Section 6, we simulate their item response to all $200$ items in the test bank, and then calculate the posterior distribution using Theorem 4.2.
Figures 10, 12, and 13 illustrate the decrease of the estimation MSE for the first quartile, median, and the third quartiles compared to those of the oracle distribution, respectively. The quartiles of each posterior distribution is computed via $1,000$ independent draws. In all these figures, we observe similar superior performances for the MI, Max Var, and the Q-learning methods over the EAP and the MAX pos approaches in estimating the entire oracle distributions. Consistent with the findings in Section 6, the Q-learning approach demonstrates the fastest error reduction rate, especially in the early stages of the test.
Figure 10: MSE Between Posterior 1st Quantile and Oracle 1st Quantile
Figure 11: MSE Between Posterior Means and Oracle Posterior Means
Figure 12: MSE Between Posterior Median and Oracle Median
Figure 13: MSE Between Posterior 3rd Quantile and Oracle 3rd Quantile
Appendix G pCAT-COG Study Additional Visualization
G.1 Exposure rate
In the real-data experiment we administer the full $57$-item bank in every simulated session to compare the rate of MSE reduction as items accrue. Consequently, item exposure is trivially $100\%$ for all items under all methods, and exposure-rate comparisons are not directly informative in this setting. Therefore, we visualize the distributions of item exposure rate when $T=30$ in Figure 14. Since the full item bank has $57$ items, we should expect the exposure rate to be centered around at around $30/57\approx 52.6\%$.
Figure 14: Distributions of Item Exposure Rates at $T=30$: pCAT-COG
G.2 Q-learning Training Dynamics
Appendix H DESE Educational Assessment
We consider an important application in educational assessment, utilizing 2022 Grade 8 student item response data from the Massachusetts Department of Elementary and Secondary Education (DESE). Excluding non-multiple-choice questions, the dataset includes responses to $24$ English items and $34$ math items. For efficiency, we randomly sampled $1,000$ student responses to evaluate each CAT item selection algorithm.
Given the unknown latent structure, we estimate the factor loading matrix via the PXL-EM algorithm [doi:10.1080/01621459.2025.2476786], resulting in a $58$ by $4$ factor loading matrix, with visualization and the details of model fitting provided in Appendix H. When estimating the item parameters, we adopted the idea of dynamic posterior exploration, letting the regularization $\lambda_{0}$ parameter in the PXL-EM algorithm to gradually increase over the path $\{1,5,10,20,40,60,80,100\}$. The resulting factor loading matrix $B$ is visualized in Figure 16, where dark black regions indicate exact zeros. Notably, all English items load exclusively onto the first factor, representing general ability, while math items can also load onto factors 2244, reflecting their greater structural complexity. For instance, the final factor (”Math 3”) has nonzero loadings only on items $34$, $36$, and $38$, all of which correspond to geometry problems.  [^4] Exam questions are available at: https://www.doe.mass.edu/mcas/2022/release/. Interestingly, the authors in [doi:10.1080/01621459.2025.2476786] obtained very similar factor loading matrix estimation for the Grade 10 data as well.
Since all items are heavily loaded on the first primary factor, representing the global latent ability, we focus on the problem of adaptively measuring the first factor while accounting for the underlying multidimensional structure in dimensions 22-44. The left subplot of Figure 15 illustrates the cumulative percentage of completed test sessions for the $1,000$ sampled examinees as more items are administered. As shown in the plot, nearly $70\%$ of CAT sessions terminate after only $20$ items under the Q-learning policy. On average over these $1,000$ examinees, Q-learning also required only $20$ items to reach the termination criterion, whereas the second-fastest method (MI) required nearly $26$ items. Beyond its efficiency in variance reduction, Q-learning also achieves a faster decay in MSE when estimating the first factor, reinforcing findings from earlier experiments.
(a) Test Termination Speed
(b) Test MSE Accuracy
Figure 15: Educational Assessment: Main Factor Posterior Variance Reduction (Left) and Estimation Accuracy (Right)
These results highlight the potential of Q-learning for educational assessments, demonstrating its ability to adapt various testing environments and to accelerate testing while maintaining accurate estimation of students’ multivariate latent traits.
Figure 16: Estimated Factor Loadings for DESE Data